### v4

# Transition Path Sampling with Improved Off-Policy Training of Diffusion Path Samplers

Abstract
Understanding transition pathways between meta-stable states in molecular systems is crucial to advance material design and drug discovery. However, unbiased molecular dynamics simulations are computationally infeasible due to the high energy barriers separating these states. Although recent machine learning techniques offer potential solutions, they are often limited to simple systems or rely on collective variables (CVs) derived from costly domain expertise. In this paper, we introduce a novel approach that trains diffusion path samplers (DPS) for transition path sampling (TPS) without the need for CVs. We recast the problem as an amortized sampling of the target path measure, minimizing the log-variance divergence between the path measure induced by our DPS and the target path measure. To ensure scalability for high-dimensional tasks, we introduce (1) a new off-policy training objective based on learning control variates with replay buffers and (2) a scale-based equivariant parameterization of the bias forces. We evaluate our approach, coined TPS-DPS, on a synthetic double-well potential and three peptides: Alanine Dipeptide, Polyproline Helix, and Chignolin. Results show that our approach produces more realistic and diverse transition pathways compared to existing baselines. We also provide links to our project page and code.
1 Introduction
In material design and drug discovery, it is crucial to understand the mechanisms and kinetics of transitions between meta-stable states of molecular systems, such as protein folding and chemical reactions (Mulholland, 2005; Piana et al., 2012; Ahn et al., 2019; Spotte-Smith et al., 2022). Their comprehensive study requires sampling transition paths (Elber, 2016; Lee et al., 2017), which provides insight into mechanisms and energy landscapes. However, naïvely sampling transition paths by molecular dynamics (MD) simulations is often computationally intractable due to high energy barriers, which cause an exponential decay in probability to make a transition (Pechukas, 1981).
To address this problem, researchers have developed enhanced sampling approaches such as steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999), umbrella sampling (Torrie & Valleau, 1977; Kästner, 2011), meta-dynamics (Ensing et al., 2006; Branduardi et al., 2012; Bussi & Branduardi, 2015), on-the-fly probability-enhanced sampling (OPES; Invernizzi & Parrinello, 2020), and adaptive biasing force (ABF; Comer et al., 2015) methods. These methods rely on bias forces to facilitate transitions across high energy barriers. They are mainly designed based on collective variables (CVs), which are functions of atomic coordinates that capture the slow modes of the transition. Although effective for some systems, the reliance on expensive domain knowledge limits the applicability of the methods to systems where CVs are less understood.
Recently, machine learning has emerged as a promising paradigm for CV-free transition path sampling (Das et al., 2021; Lelièvre et al., 2023; Holdijk et al., 2024). The key idea is to parameterize the bias force using a neural network and train the corresponding biased MD simulation to sample realistic transition paths. In particular, Lelièvre et al. (2023) considered reinforcement learning to sample paths escaping meta-stable states. Das et al. (2021) and Holdijk et al. (2024) proposed to minimize the Kullback-Leibler (KL) divergence between the path measures induced by the neural network and the target path measure. However, minimizing the KL divergence leads to mode collapse, which is hazardous for sampling transition paths due to a lack of exploration. Furthermore, Das et al. (2021) and Lelièvre et al. (2023) limited their evaluation to low-dimensional synthetic systems. Designing machine learning algorithms for CV-free transition path sampling of large molecules remains an open problem.
Contribution. In this work, we propose to solve the transition path sampling problem with the diffusion path sampler (TPS-DPS). TPS-DPS trains the bias force to minimize a recently proposed log-variance divergence (Nüsken & Richter, 2021) between the path measure induced by the bias MD and the target path measure. To ensure scalability for high-dimensional tasks, e.g., Chignolin folding, we propose a new off-policy training scheme and a scale-based parameterization of the bias forces.
To be specific, we propose a new training objective to minimize the log-variance divergence, defined between a pair of path measures induced by the biased MD and the unbiased MD conditioned on the target meta-stable state. To minimize the variance of gradient estimators and improve sample efficiency, we propose to use (1) learnable control variates and (2) replay buffers with simulated annealing. Our choice of minimizing the log-variance is motivated by recent works (Richter et al., 2020; Nüsken & Richter, 2021; Richter & Berner, 2024) that demonstrate many desirable properties of the log-variance divergence such as a robust gradient estimator for training stability and avoidance of differentiation through the SDE solver. This allows the use of previously collected samples in a replay buffer, promoting sample efficiency in training and improving the diversity of the generated transition paths by preventing the mode collapse observed in KL divergence.
We also introduce a new $SE(3)$ equivariant scale-based parameterization for the bias force that guides the biased MD to sample meaningful paths more frequently, based on inductive bias. Our key idea is to parameterize the bias force to yield a positive inner product with a direction to decrease the distance between the molecular state and the target meta-stable state, which guarantees the bias force to (locally) decrease the distance. We also use the Kabsch algorithm (Kabsch, 1976) to align the molecule and the bias force to the target meta-stable state, guaranteeing $SE(3)$ equivariance for better generalization across molecular states that are identical up to roto-translations.
We extensively evaluate our method on the synthetic double-well potential and three peptides: Alanine Dipeptide, Polyproline Helix, and Chignolin. We compare TPS-DPS with prior ML approach (PIPS; Holdijk et al., 2024), as well as classical non-ML methods, e.g., steered MD (SMD; Izrailev et al., 1999). Our experiments demonstrate that TPS-DPS consistently generates realistic and diverse transition paths, similar to the ground truth ensemble. In addition, ablation studies of algorithmic components verify the effectiveness of our method.
2 Related work
Transition path sampling without ML. Metadynamics (Branduardi et al., 2012), on-the-fly probability-enhanced sampling (OPES; Invernizzi & Parrinello, 2020), adaptive biasing force (ABF; Comer et al., 2015), and steered molecular dynamics (SMD; Schlitter et al., 1994; Izrailev et al., 1999) were introduced to explore molecular conformations that are difficult to access by unbiased molecular dynamics (MD) within limited simulation times (Hénin et al., 2022). However, they mostly rely on collective variables (CVs) for high-dimensional problems and are inapplicable to systems with unknown CVs. To sample transition paths without CVs, Dellago et al. (1998) proposed shooting methods that use the Markov chain Monte Carlo (MCMC) procedure on path space.
Data driven ML approaches. Recently, generative models have been trained to sample a new transition path given a dataset of transition paths. Petersen et al. (2023); Triplett & Lu (2023) and Lelièvre et al. (2023) applied diffusion probabilistic models (Ho et al., 2020) and variational auto-encoders (Kingma & Welling, 2013) for transition path sampling, respectively. However, these methods are limited to small systems. Concurrent to our work, Jing et al. (2024) generated molecular trajectories by task-wise conditions given an initial state and viewed transition path sampling as an interpolation task between two molecular states.
We also note some related works that are not directly comparable to our method. Klein et al. (2024) proposed to accelerate MD simulation through coarse-grained path generation that may include multiple meta-stable states. However, this work is incapable of targeting a specific meta-stable state and generating fine-grained transition paths. Duan et al. (2023); Kim et al. (2024) use neural networks to generate transition states of a given chemical reaction, but cannot generate transition paths.
Data free ML approaches. Without a previously collected dataset, Das et al. (2021); Lelièvre et al. (2023); Hua et al. (2024); Holdijk et al. (2024) trained generators to sample transition paths using biased MD. Lelièvre et al. (2023) used reinforcement learning to train the generator but focused on escaping an initial meta-stable state rather than reaching the target meta-stable state. Next, Holdijk et al. (2024) considered transition path sampling as minimizing the reverse KL divergence between path distribution from biased MD and transition path distribution. In this work, we mainly compare our method with Holdijk et al. (2024) and provide quantitative analysis for real-world molecule systems. We also note that, concurrent to our work, Du et al. (2024) also considered variational formulation of the TPS problem using Doob’s $h$-transform.
3 Transition path sampling with diffusion path samplers
Figure 1: Problem setup. Potential energy landscape of Alanine Dipeptide, where the sampled path from state $\mathcal{A}$ to state $\mathcal{B}$ is highlighted in yellow dotted lines. We visualize snapshots in the transition path highlighted in white circles and the transition state highlighted in the white star.
In this section, we introduce our method, coined transition path sampling with diffusion path sampling (TPS-DPS). Our main idea is to formulate the transition path sampling (TPS) problem as a minimization of log-variance divergence (Nüsken & Richter, 2021) between two path measures: the path measure of transition paths and that of biased MD simulations parameterized by a neural network. Our main methodological contribution is twofold: (1) a new training algorithm that minimizes the log-variance divergence with a learned control variate and replay buffer, and (2) a $SE(3)$ equivariant scale-based parameterization of the biased MD that provides inductive bias for dense training signals in high-dimensional problems.
3.1 Problem setup
Our goal is to sample transition paths from one meta-stable state to another meta-stable state given a molecule system. We provide an example of the problem for Alanine Dipeptide in Figure 1. We view this as a task to sample paths from an unbiased molecular dynamics (MD), described by a stochastic differential equation (SDE), conditioned on its starting and ending points. To solve this task, we train the bias force parameterized by a neural network to amortize the sampling procedure.
Molecular dynamics. We consider a MD simulation on time interval $[0,T]$ that describes motion of a molecular state $\bm{X}_{t}=(\bm{R}_{t},\bm{V}_{t})\in\mathbb{R}^{6N}$ at time $t$ where $N$ is the number of atoms, $\bm{R}_{t}\in\mathbb{R}^{3N}$ is the atom-wise positions and $\bm{V}_{t}\in\mathbb{R}^{3N}$ is the atom-wise velocities. In particular, we adopt Langevin dynamics (Bussi & Parrinello, 2007) defined as the following SDE:
$\mathrm{d}\bm{X}_{t}=\bm{u}(\bm{X}_{t})\mathrm{d}t+\Sigma\mathrm{d}\bm{W}_{t},~{}~{}\bm{u}(\bm{X}_{t})=\left(\bm{V}_{t},-\frac{\nabla U(\bm{R}_{t})}{\bm{m}}-\gamma\bm{V}_{t}\right),~{}~{}\Sigma=\text{diag}\left(\bm{\zeta},\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\right)$ (1)
where $U$, $\bm{m}$, $\gamma$, $k_{B}$, $\lambda$, and $\bm{W}_{t}$ denote the potential energy function, the atom-wise masses, the friction term, the Boltzmann constant, the absolute temperature, and the Brownian motion, respectively, and $\bm{\zeta}\in\mathbb{R}^{3N}$ is a vector of positive infinitesimal values. We let $\mathbb{P}_{0}$ denote the path measure induced by Equation 1 in the space $\mathcal{C}([0,T];\mathbb{R}^{6N})$ of continuous functions $\bm{X}:[0,T]\rightarrow\mathbb{R}^{6N}$.
Transition path sampling. One of the challenges in sampling transition paths through unbiased MD simulations is the meta-stability: a state remains trapped for a long time in the initial meta-stable state $\mathcal{A}\subseteq\mathbb{R}^{3N}$ before transitioning into a distinct meta-stable state $\mathcal{B}\subseteq\mathbb{R}^{3N}$. To capture the rare event where transition from $\mathcal{A}$ to $\mathcal{B}$ occurs, we constrain paths $\bm{X}=(\bm{X}_{t})_{0\leq t\leq T}$ sampled from unbiased MD to satisfy $\bm{R}_{0}\in\mathcal{A}$, $\bm{R}_{T}\in\mathcal{B}$ for a fixed time $T$. Since the meta-stable state $\mathcal{A}$ and $\mathcal{B}$ are not well-specified for many molecular systems, we simplify this task by (1) fixing a local minima $\bm{R}_{\mathcal{A}},\bm{R}_{\mathcal{B}}$ of the potential energy function in the meta-stable states $\mathcal{A},\mathcal{B}$ and (2) sampling a transition path $\bm{X}$ that starts from the state $\bm{R}_{0}=\bm{R}_{\mathcal{A}}$ and ends at the vicinity of $\bm{R}_{\mathcal{B}}$.
To be specific, we aim to sample from the target path measure $\mathbb{Q}$ which reweights the path measure $\mathbb{P}_{0}$ via the Radon–Nikodym derivative, defined as follows:
$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\frac{1_{\mathcal{B}}(\bm{X})}{\mathbb{E}_{\mathbb{P}_{0}}\left[1_{\mathcal{B}}(\bm{X})\right]},\quad 1_{\mathcal{B}}(\bm{X})=\begin{cases}1&\text{if}~{}\lVert\bm{R}_{T}-\rho_{T}\cdot\bm{R}_{\mathcal{B}}\rVert\leq\delta,\\ 0&\text{otherwise},\end{cases}$ (2)
where $\cdot$ denotes group action associated with the $SE(3)$ space and $\rho_{T}\cdot\bm{R}_{\mathcal{B}}$ is the aligned position by the optimal roto-translation $\rho_{T}\in SE(3)$ to minimize its Euclidean distance to $\bm{R}_{T}$, i.e., $\rho_{T}=\operatorname*{arg\,min}_{\rho\in SE(3)}\lVert\bm{R}_{T}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$. Such a transformation can be obtained from the Kabsch algorithm in $O(N)$ complexity (Kabsch, 1976).
Note that one may consider naïve rejection sampling to sample transition paths, based on running unbiased MD to sample a path $\bm{X}$ from the path measure $\mathbb{P}_{0}$ and accepting if the path $\bm{X}$ arrives at the neighborhood of the position $\bm{R}_{\mathcal{B}}$ with the radius $\delta$. However, this method does not scale to high-dimensional or low-temperature problems, since the sampled path by unbiased MD rarely reaches the target states due to the high energy barriers, i.e., the rejection ratio is too high.
3.2 Log-variance minimization
In this section, we propose our algorithm to amortize transition path sampling. Our key idea is to train a neural network to induce a path measure that matches the target path measure $\mathbb{Q}$, using the log-variance divergence (Nüsken & Richter, 2021) between the path measures. We propose a new training scheme to minimize the log-variance divergence based on learning the control variate of its gradient and a replay buffer to improve sample efficiency and diversity.
Amortizing transition path sampling with log-variance divergence. To match the target path measure $\mathbb{Q}$, we consider a biased MD defined by a policy $\bm{v}$ (or bias force $\bm{b}$) as the following SDE:
$\mathrm{d}\bm{X}_{t}=(\bm{u}(\bm{X}_{t})+\Sigma\bm{v}(\bm{X}_{t}))\mathrm{d}t+\Sigma\mathrm{d}\bm{W}_{t},\quad\bm{v}(\bm{X}_{t})=\Sigma^{-1}\left(\bm{0},\frac{\bm{b}(\bm{X}_{t})}{\bm{m}}\right).$ (3)
We also let $\mathbb{P}_{\bm{v}}$ denote the path measure induced by the SDE. To amortize transition path sampling, we match the path measure $\mathbb{P}_{\bm{v}_{\theta}}$ of a parameterized policy $\bm{v}_{\theta}$ with the target path measure $\mathbb{Q}$ by minimizing the log-variance divergence:
$D_{\text{LV}}^{\mathbb{P}}(\mathbb{P}_{\bm{v}_{\theta}}\|\mathbb{Q})=\mathbb{V}_{\mathbb{P}}\left[\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}\right]=\mathbb{E}_{\mathbb{P}}\left[\left(\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}-\mathbb{E}_{\mathbb{P}}\left[\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}\right]\right)^{2}\right],$ (4)
where $\mathbb{P}$ is an arbitrary reference path measure with $\mathbb{E}_{\mathbb{P}}[\log({\mathrm{d}\mathbb{Q}}/{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}})]<\infty$. To express the log-variance divergence in detail, we let $\mathbb{P}=\mathbb{P}_{\tilde{\bm{v}}}$ for some policy $\tilde{\bm{v}}$ and apply the Girsanov’s theorem to Equation 4, deriving the following formulation:
$D_{\text{LV}}^{\mathbb{P}_{\tilde{\bm{v}}}}(\mathbb{P}_{\bm{v}_{\theta}}\|\mathbb{Q})=\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}\left[(F_{\bm{v}_{\theta},\tilde{\bm{v}}}-\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}[F_{\bm{v}_{\theta},\tilde{\bm{v}}}])^{2}\right]\,,$ (5)
$F_{\bm{v}_{\theta},\tilde{\bm{v}}}(\bm{X})=\frac{1}{2}\int^{T}_{0}\lVert\bm{v}_{\theta}(\bm{X}_{t})\rVert^{2}\mathrm{d}t-\int^{T}_{0}(\bm{v}_{\theta}\cdot\tilde{\bm{v}})(\bm{X}_{t})\mathrm{d}t-\int^{T}_{0}\bm{v}_{\theta}(\bm{X}_{t})\cdot\mathrm{d}\bm{W}_{t}+\log 1_{\mathcal{B}}(\bm{X}).$ (6)
The first three terms in Equation 6 correspond to the deviation of the biased MD from the unbiased MD integrated over the path sampled from $\mathbb{P}_{\tilde{\bm{v}}}$. The last term reweights the unbiased MD to the target path measure $\mathbb{Q}$. As a result, minimizing Equation 5 could be thought as minimizing the variation between $\mathbb{P}_{\bm{v}_{\theta}}$ and $\mathbb{Q}$. We provide the full derivation in Section A.1. Compared to KL divergence, the log-variance divergence provides a robust gradient estimator and avoids differentiating through the SDE solver. (Richter et al., 2020; Nüsken & Richter, 2021).
Minimizing with learnable control variate. To minimize the log-variance divergence, we consider the following loss function that replaces the estimation of $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ by learning a parameter $w$:
$\mathcal{L}(\theta,w)=\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}\left[(F_{\bm{v}_{\theta},\bm{v}_{\theta}}-w)^{2}\right],$ (7)
where $w$ is a control variate that controls the variance of the gradient estimator of $\nabla_{\theta}\mathcal{L}(\theta,w)$ without changing the gradient. Note that we set $\tilde{\bm{v}}=\bm{v}_{\theta}$ in Equation 6, which implies that the gradient of Equation 7 coincides with the KL divergence (Richter et al., 2020; Nüsken & Richter, 2021). When optimized, the control variate $w$ estimates the expectation $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ since $\operatornamewithlimits{argmin}_{w}\mathcal{L}(\theta,w)=\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$. Thus, jointly optimizing $(\theta,w)$ with the gradient step can be interpreted as jointly minimizing log-variance divergence and estimating $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ using $w$.
Off-policy training with replay buffer. To leverage the degree of freedom in reference path measure for the log-variance divergence, we allow discrepancy between reference path measure and current path measure, called off-policy training, which is widely used in discrete-time reinforcement learning (Mnih et al., 2013; Bengio et al., 2021). For the sample efficiency, we reuse the samples with a replay buffer $\mathcal{D}$ which stores path samples from the path measure $\mathbb{P}_{\bm{v}_{\bar{\theta}}}$ associated with previous policies $\bm{v}_{\bar{\theta}}$. Our modified loss function $\mathcal{L}^{\mathcal{D}}$ with replay buffer $\mathcal{D}$ is defined as follows:
$\mathcal{L}^{\mathcal{D}}(\theta,w)=\mathbb{E}_{(\bm{v}_{\bar{\theta}},\bm{X})\sim\mathcal{D}}[(F_{\bm{v}_{\theta},\bm{v}_{\bar{\theta}}}(\bm{X})-w)^{2}].$ (8)
Using the replay buffer also prevents mode collapse, using diverse paths from different path measures. Similar to other off-policy training algorithms (Malkin et al., 2022; Kim et al., 2023), we use simulated annealing to collect diverse paths that cross high energy barriers.
Discretization. To implement the algorithm, we discretize Equation 8. Given a discretization step size $\Delta t$, we consider the discretized paths $\bm{x}_{0:L}=(\bm{x}_{0},\bm{x}_{1},\ldots,\bm{x}_{L})$ of $\bm{X}$ from MD simulations where $L=T/\Delta t$ and $\bm{x}_{\ell}=\bm{X}(\ell\Delta t)$. In discrete cases, the discretized paths $\bm{x}_{0:L}$ from previous policies $\bm{v}_{\bar{\theta}}$ and their (gradient-detached) policy values $(\bm{v}_{\bar{\theta}}(\bm{x}_{0}),...,\bm{v}_{\bar{\theta}}(\bm{x}_{L}))$ are used to approximate the value $F_{\bm{v}_{\theta},\bm{v}_{\bar{\theta}}}(\bm{X})$ in Equation 6 as follows:
$\hat{F}_{\bm{v}_{\theta},\bm{v}_{\bar{\theta}}}(\bm{x}_{0:L})=\frac{1}{2}\sum_{\ell=0}^{L-1}\lVert\bm{v}_{\theta}(\bm{x}_{\ell})\rVert^{2}\Delta t-\sum_{\ell=0}^{L-1}(\bm{v}_{\theta}\cdot\bm{v}_{\bar{\theta}})(\bm{x}_{\ell})\Delta t-\sum_{\ell=0}^{L-1}\bm{v}_{\theta}(\bm{x}_{\ell})\cdot\bm{{\epsilon}}_{\ell}+\log 1_{\mathcal{B}}(\bm{x}_{0:L})\,,$ (9)
where the noise $\bm{{\epsilon}}_{\ell}=\Sigma^{-1}(\bm{x}_{\ell+1}-\bm{x}_{\ell}-(\bm{u}(\bm{x}_{\ell})+\Sigma\bm{v}_{\bar{\theta}}(\bm{x}_{\ell}))\Delta t)$ is the discretized Brownian motions of the Langevin dynamics with policy $\bm{v}_{\bar{\theta}}$. For implementation, we further derive a simple discretized loss of Equation 8 from Equation 9 as follows:
$\mathbb{E}_{\bm{x}_{0:L}\sim\hat{\mathcal{D}}}\left[\left(\log\frac{p_{0}(\bm{x}_{0:L})1_{\mathcal{B}}(\bm{x}_{0:L})}{p_{\bm{v}_{\theta}}(\bm{x}_{0:L})}-w\right)^{2}\right],$ (10)
where the buffer $\hat{\mathcal{D}}$ stores paths $\bm{x}_{0:L}$ sampled from the previous policies, and $p_{0}$ and $p_{\bm{v}_{\theta}}$ denote discrete time transition probability induced by Equations 1 and 3, respectively. We provide a formal derivation of the discretized loss in Section A.2. Note that the same objective was derived in the name of relative trajectory balance by Venkatraman et al. (2024).
We describe our training algorithm in Section A.3. Overall, our off-policy training algorithm iterates through four steps: (1) sampling paths from the biased MD simulation with current policy $\bm{v}_{\theta}$ at high temperature, (2) storing sampled paths in the replay buffer $\hat{\mathcal{D}}$, (3) sampling a batch of the paths from the replay buffer, and (4) training current policy $\bm{v}_{\theta}$ by minimizing the loss in Equation 10. After minimization, biased MD simulation can directly sample transition paths by amortizing inference in the target path measure.
3.3 Parameterization for high dimensional tasks
In this section, we introduce new parameterizations of the bias force and the indicator function for high-dimensional tasks. Our parameterization is designed around alleviating the problem of sparse training signal, where the model is unable to learn anything due to the failure of collecting meaningful paths that end near the meta-stable state. This problem is especially severe in high-dimensional problems such as large molecules.
Bias force parameterization. To frequently sample the meaningful paths, we parameterize the bias force to guarantee (locally) decreasing the distance between the current molecular state and the target meta-stable state. This is done by parameterizing the bias force as positive atom-wise scaling of the force minimizing the distance, which implies that the bias force can be decomposed into two terms: (1) the direction to decrease the Euclidean distance and (2) the direction that leaves the Euclidean distance invariant. The bias force also satisfies roto-translational equivariance to the current state $\bm{X}_{t}$, aligning with the symmetry of the transition path sampling problem for better generalization.
To be specific, we use a roto-translation $\rho_{t}\in SE(3)$ to align $\bm{R}_{\mathcal{B}}$ with $\bm{R}_{t}$, as we do for indicator function in Equation 2. To achieve $SE(3)$ equivariant, we parameterize the bias force as follows:
$\bm{b}(\bm{X}_{t})=\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))(\rho_{t}\cdot{\bm{R}}_{\mathcal{B}}-\bm{R}_{t}),$ (11)
where $\bm{s}_{\theta}(\cdot)\in\mathbb{R}_{+}^{3N}$ is a neural network constrained to have positive output elements. One can observe the positive inner product between the bias force and the direction to minimize the Euclidean distance, i.e., $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$, as follows:
$\bm{b}(\bm{X}_{t})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})=(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})^{T}\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})\geq 0,$ (12)
which implies that the bias force can be decomposed into a vector positively aligning with $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$ and a vector orthogonal to $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$.
We also note that the bias force is $SE(3)$ equivariant to the current state $\bm{X}_{t}$ since it is the product of an invariant term $\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))$ and an equivariant term $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$. This stems from the fact that the Kabsch algorithm is equivariant to the current state $\bm{X}_{t}$, i.e., roto-translating position $\bm{R}_{t}\mapsto\bar{\rho}\cdot\bm{R}_{t}$ leads to roto-translating output of the Kabsch algorithm $\rho_{t}\mapsto\bar{\rho}\cdot\rho_{t}$.
In the experiments, we also consider other equivariant parameterizations that are less constrained: (1) directly predicting the bias force by $\rho_{t}\cdot\bm{b}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}^{3N}$ and (2) predicting the force as gradient of the invariant bias potential $\nabla b_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}$. We observe the parameterizations to be useful for low-dimensional tasks, but suffer from the sparse training signal on high-dimensional tasks.
Indicator function parameterization. We also propose to relax the indicator function $1_{\mathcal{B}}$ as a radial basis function (RBF) kernel $\tilde{1}_{\mathcal{B}}(\bm{X})=k(\bm{R}_{T},\rho^{-1}\cdot{\bm{R}}_{\mathcal{B}};\sigma^{2})$ which measures the similarity between two positions and $\sigma>0$ controls the degree of relaxation. The range of RBF kernel $k$ is bounded by the interval $(0,1]$ so that $\log\tilde{1}_{\mathcal{B}}(\bm{X})$ is well-defined and $\tilde{1}_{\mathcal{B}}(\bm{X})$ represents the binary indicator function smoothly. To capture high training signal, we focus on the signal from subtrajectories of sampled paths, taking maximum over RBF kernel values of all intermediate states by $\tilde{1}_{\mathcal{B}}^{\text{max}}(\bm{X})=\max_{t\in[0,T]}k(\bm{R}_{t},\tilde{\bm{R}}_{\mathcal{B}};\sigma^{2})$. To extract the subtrajectory with a high training signal, we can truncate the paths at the time that maximizes RBF kernel values, allowing variable path lengths. Also, the relaxed indicator function is $SE(3)$ invariant to $\bm{R}_{t}$ due to the Kabsch algorithm.
4 Experiment
Table 1: Benchmark scores on the double-well system and three real-world molecular systems: Alanine Dipeptide, Polyproline Helix, and Chignolin. All metrics are averaged over 1024 paths for the double-well system, and 64 paths for real-world molecular systems. ETS is computed for paths that hit the target meta-stable state, and the best results are highlighted in bold. TPS-DPS predicting the bias force, potential, and atom-wise scaling are denoted by (F), (P), and (S), respectively. UMD ($\lambda$) denotes unbiased MD with temperature $\lambda$ and SMD ($k$) denotes steered MD with the force constant $k$. Unless otherwise specified, paths are generated by MD simulation at 1200K for double-well and 300K for real-world molecules.
Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
Å % $\mathrm{kJmol}^{-1}$
Double-well
UMD (1200K) 2.21 $\pm$ 0.10 0.00 -
UMD (2400K) 2.11 $\pm$ 0.38 3.03 1.69 $\pm$ 0.31
UMD (3600K) 1.85 $\pm$ 0.68 12.60 2.12 $\pm$ 0.41
UMD (4800K) 1.54 $\pm$ 0.81 21.58 2.77 $\pm$ 0.69
SMD (0.5) 0.98 $\pm$ 0.90 52.15 1.54 $\pm$ 0.21
SMD (1) 0.14 $\pm$ 0.08 99.80 1.85 $\pm$ 0.16
TPS-DPS (F, Ours) 0.01 $\pm$ 0.02 99.90 1.38 $\pm$ 0.16
TPS-DPS (P, Ours) 0.01 $\pm$ 0.03 99.71 1.36 $\pm$ 0.15
TPS-DPS (S, Ours) 0.01 $\pm$ 0.03 99.80 1.73 $\pm$ 0.20
Polyproline Helix
UMD (300K) 2.22 $\pm$ 0.11 0.00 -
UMD (1200K) 1.38 $\pm$ 0.45 10.94 1010.28 $\pm$ 38.44
SMD (5k) 1.68 $\pm$ 0.17 54.69 350.58 $\pm$ 14.36
SMD (10k) 1.26 $\pm$ 0.06 100.00 355.62 $\pm$ 14.83
PIPS (F) 2.64 $\pm$ 0.15 0.00 -
PIPS (P) 1.85 $\pm$ 0.13 93.75 574.66 $\pm$ 20.49
TPS-DPS (F, Ours) 1.53 $\pm$ 0.12 98.44 418.17 $\pm$ 45.54
TPS-DPS (P, Ours) 1.35 $\pm$ 0.12 100.00 345.00 $\pm$ 32.58
TPS-DPS (S, Ours) 1.17 $\pm$ 0.02 100.00 342.00 $\pm$ 20.28
Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
Å % $\mathrm{kJmol}^{-1}$
Alanine Dipeptide
UMD (300K) 1.59 $\pm$ 0.15 0.00 -
UMD (3600K) 1.19 $\pm$ 0.32 6.25 812.47 $\pm$ 148.80
SMD (10) 0.86 $\pm$ 0.21 7.81 33.15 $\pm$ 6.46
SMD (20) 0.56 $\pm$ 0.27 54.69 78.40 $\pm$ 12.76
PIPS (F) 0.66 $\pm$ 0.15 43.75 28.17 $\pm$ 10.86
PIPS (P) 1.66 $\pm$ 0.03 0.00 -
TPS-DPS (F, Ours) 0.16 $\pm$ 0.06 92.19 19.82 $\pm$ 15.88
TPS-DPS (P, Ours) 0.16 $\pm$ 0.10 87.50 18.37 $\pm$ 10.86
TPS-DPS (S, Ours) 0.25 $\pm$ 0.20 76.00 22.79 $\pm$ 13.57
Chignolin
UMD (300K) 7.98 $\pm$ 0.41 0.00 -
UMD (1200K) 7.23 $\pm$ 0.93 1.56 388.17
SMD (10k) 1.26 $\pm$ 0.31 6.25 -527.95 $\pm$ 93.58
SMD (15k) 1.17 $\pm$ 0.31 23.44 -237.15 $\pm$ 122.29
PIPS (F) 4.66 $\pm$ 0.17 0.00 -
PIPS (P) 4.67 $\pm$ 0.32 0.00 -
TPS-DPS (F, Ours) 4.41 $\pm$ 0.49 0.00 -
TPS-DPS (P, Ours) 3.87 $\pm$ 0.42 0.00 -
TPS-DPS (S, Ours) 1.17 $\pm$ 0.66 59.38 -780.18 $\pm$ 216.93
In this section, we compare our method, called TPS-DPS, with both classical and ML approaches, assessing the accuracy and diversity of sampled transition paths. We begin with a synthetic double-well system at 1200K, followed by three real-world molecular systems with various numbers of amino acids: Alanine Dipeptide, Polyproline Helix, and Chignolin. Additionally, we conduct ablation studies to validate the effectiveness of each component in our method. All real-world molecular systems are simulated using the OpenMM library (Eastman et al., 2023). Details on OpenMM simulation and model configurations are provided in Sections B.1 and B.2, respectively.
Evaluation Metrics. We consider three metrics for evaluation: RMSD, THP, and ETS. The root mean square distance (RMSD) calculates the Euclidean distance between the atoms in the final position of the sampled path and the target position $\bm{R}_{\mathcal{B}}$, with the final position aligned to the target. The target hit percentage (THP) measures the number of paths that successfully arrive at the target meta-stable state $\mathcal{B}$. Finally, the energy of the transition state (ETS) measures the ability to identify probable transition states. For further details, refer to Section B.3.
Baselines. We compare TPS-DPS with both non-ML and ML baselines. For non-ML baselines, we consider unbiased MD (UMD) in various temperatures and steered MD (SMD; Izrailev et al., 1999) with various force constants $k$ and collective variables. For ML baselines, we consider a CV-free transition path sampling method, path integral path sampling (PIPS; Holdijk et al., 2024) which also trains a bias force by minimizing the KL divergence between path measures induced by the biased MD and the target path measure. For simplicity, we denote parameterizations for predicting force, potential, and atom-wise scaling as (F), (P) and (S), respectively.
(a) Double-well energy landscapes
(b) Distributions of the energy and $y$
Figure 2: Visualization of potential energy landscapes and distributions of the double-well system. (2(a)) Visualization of the neural bias potential $b_{\theta}$ for TPS-DPS (P) in the double-well potential energy landscape. (2(b)) Distributions of the potential energy (left) and $y$ coordinate (right) of transition states from 1024 transition paths sampled by each method.
4.1 Double-well system
(a) UMD
(b) SMD ($0.5$)
(c) SMD ($1$)
(d) TPS-DPS (F)
(e) TPS-DPS (P)
(f) Ground Truth
Figure 3: 16 transition path sampled from each method on the potential energy landscape of the double-well system. White circles indicate the meta-stable states and white stars indicate the saddle points. We sample paths from the left meta-stable state to the right. SMD ($k$) denotes steered MD with the force constant $k$. All paths are generated by MD simulation at 300K. SMD ($k$) denotes steered MD with the force constant $k$.
We begin by evaluating our method on a two-dimensional synthetic system, i.e., a double-well potential at 1200K. This system has two global minima (white circles) representing the meta-stable states, and two reaction pathways via saddle points (white stars). We sample transition paths from the left meta-stable state $\bm{R}_{\mathcal{A}}$ to the right meta-stable states =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\bm{R}-\bm{R}_{\mathcal{B}}\rVert<0.5\}$, using over-damped Langevin dynamics. We collect ground truth path ensembles by rejection sampling which proposes paths sampled from the unbiased MD simulations and accepts if the final states are in the target meta-stable states $\mathcal{B}$. We provide more details on the system in Section B.4.
In Table 1, TPS-DPS shows superior performance for the double-well potential regardless of the bias force design. In Figure 2, the neural bias potential accelerates the transition by increasing the potential energy near the initial meta-stable state while decreasing the potential energy near the two energy barriers. Moreover, the distribution of energy and $y$ coordinates of the transition states from TPS-DPS is closest to the ground truth compared with other baselines, successfully capturing two reaction channels. In Figure 3, unbiased MD at 1200K fails to escape the initial state while SMD struggles to pass the saddle points. Our approach generates transition paths more similar to the ground truth than other methods.
(a) UMD
(b) SMD (10)
(c) SMD (20)
(d) PIPS (F)
(e) TPS-DPS (F)
(f) TPS-DPS (P)
Figure 4: 16 sampled paths from each method on the Ramachandran plot of Alanine Dipeptide. White circles indicate meta-stable states and white stars indicate saddle points. We sample transition paths from the meta-stable state $C5$ (upper left) to $C7ax$ (lower right). SMD ($k$) denotes steered MD with the force constant $k$. Paths are generated by MD simulation at 1200K.
Figure 5: Snapshots of transition path for Alanine Dipeptide (left) and Polyproline Helix (right).
(a) UMD
(b) SMD (10k)
(c) SMD (15k)
(d) PIPS (P)
(e) TPS-DPS (P)
(f) TPS-DPS (S)
Figure 6: 16 sampled paths from each method projected to the top two TICA components. White circles indicate meta-stable states. We sample transition paths from the unfolded state (right) to the folded state (left). SMD ($k$) denotes steered MD with the force constant $k$. All paths are generated by MD simulation at 300K. (P) and (S) refer to predicting the bias potential and atom-wise scaling.
4.2 Alanine Dipeptide conformation change
We first consider Alanine Dipeptide for a real-world molecule consisting of two alanine residues, sampling transition paths from the $C5$ (upper left) to the $C7ax$ (lower right) as seen in Figure 4. The target meta-stable states are defined as =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$, where $\xi(\bm{R})=(\phi,\psi)$ is a well-known collective variable which consists of two backbone dihedral angles. Also, Alanine Dipeptide has two reaction channels between the $C5$ and $C7ax$ state passing through the saddle points marked as white stars.
In Table 1 and Figure 4, unbiased MD at 300K fails to escape the initial state, while SMD with the collective variable $\xi$ generates transition paths with less probable transition states. PIPS generates transition paths of only one reaction channel, suffering from mode collapse. Compared to the baselines, our method successfully generates diverse transition paths that pass two reaction channels and show superior performance regardless of the bias force design. For qualitative analysis, we plot the snapshots of the transition path sample from TPS-DPS in Figure 5.
4.3 Polyproline Helix isomerization
Next, we consider a more complex molecule, Polyproline Helix consisting of three proline residues. We sample transition paths from the left-handed state (PPII) in the cis-configuration to the right-handed state (PPI) in the trans-configuration. For simplicity, we define the target meta-stable state based on handedness $H$ as =$\mathcal{B}=\{\bm{R}~{}|~{}H(\bm{R})>0\}$. For the formal definition of handedness, refer to Section B.4. As seen in Table 1, our models consistently outperform all baselines: unbiased MD, SMD, and PIPS. In Figure 7(a), a transition path sampled from TPS-DPS crosses the energy barrier around 0.4ps and changes handedness for 1.5ps. Additionally in Figure 5, we plot the snapshots of the sampled transition path.
(a) Potential energy and CVs over time.
(b) Visualization of hydrogen bond formation in Chignolin.
Figure 7: Qualitative evaluation on transition path sampled from TPS-DPS. (7(a)) Potential energies and collective variables (CVs) of Alanine dipeptide, Polyporline Helix, and Chignolin from top to bottom. For CVs, we plot the distance of the backbone dihedral angles between the current state and the target state for Alanine Dipeptide, the handedness of the backbone for Polyproline Helix, and the hydrogen bond distance for Chignolin. One can see that the sampled path successfully transitioned to the target meta-stable state. (7(b)) Visualization of hydrogen bond formation for Chignolin in a sampled path. We project a sampled path on the top two TICA components. We visualize the time steps indicated by the black circles, with each hydrogen bond highlighted in green and yellow.
4.4 Chignolin folding
Finally, we consider a challenging molecule, Chignolin, an artificial protein consisting of 10 amino acids (Honda et al., 2004), which folds into a $\beta$-hairpin structure by hydrogen bonds. We sample transition paths from the unfolded state (right white circle) to the folded state (left white circle) in Figure 6. We define the target meta-stable state =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$ where $\xi$ consists of the top two components of time-lagged independent component analysis (TICA; Pérez-Hernández et al., 2013). We further describe TICA in Section B.4.
[…]
4.5 Ablation study
Effectiveness of algorithmic components. First, we conduct ablation studies to verify the effectiveness of the five proposed components: log-variance loss, learnable control variate, replay buffer, simulated annealing, and maximum over RBF values for various path lengths. To be specific, we (1) replace our loss with the KL divergence, (2) replace the learnable control variate with the local control variate used in Nüsken & Richter (2021), (3) remove the replay buffer and use data only from the current policy, (4) use only one temperature $\lambda$, and (5) remove maximum operation over RBF values using only the final state.
[…]
As seen in Figure 9, all the proposed components improve performance. Our loss is smaller than the KL divergence by more than two orders of magnitude and significantly improves performance. Learning the control variate slightly improves performance, showing that utilizing data from previous policies is effective. The replay buffer significantly improves training efficiency, and shows that the large performance gap between our loss and KL divergence comes from the replay buffer. Simulated annealing for biased MD simulation is critical to finding transition paths. RMSD does not decrease without simulated annealing while loss decreases significantly. For the relaxed indicator function, maximum operation accelerates convergence and improves performance with frequent training signals from the subtrajectories.
Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
Å % $\mathrm{kJmol}^{-1}$
Ours 0.16 $\pm$ 0.06 92.19 19.82 $\pm$ 15.88
w/ KL 0.43 $\pm$ 0.34 53.12 27.88 $\pm$ 14.38
w/ local 0.24 $\pm$ 0.15 73.44 22.53 $\pm$ 14.45
w/o replay 0.33 $\pm$ 0.27 64.06 24.38 $\pm$ 12.31
w/o annealing 0.67 $\pm$ 0.21 9.38 69.86 $\pm$ 30.15
w/o various len 0.23 $\pm$ 0.11 75.00 29.49 $\pm$ 14.13
(a) Component-wise performance on Alanine Dipeptide.
(b) Loss and RMSD curve over rollouts.
Figure 9: Ablation studies on the components of TPS-DPS in the Alanine Dipeptide. (9(a)) Benchmark scores on Alanine Dipeptide. All metrics are averaged over 64 paths, and we highlight the best results in bold. ETS is computed for paths that hit the target meta-stable state. (9(b)) Loss and RMSD curves averaged over 8 different seeds.
Comparison to steered MD with varying force constant $k$. The performance of SMD differs by the force constant $k$. As seen in Figure 8, one can see the trade-off between THP and ETS for various $k$ values. TPS-DPS outperforms SMD in THP and ETS without showing a trade-off between them.
5 Conclusion
In this work, we introduced a novel CV-free diffusion path sampler, called TPS-DPS, to amortize transition path sampling, using log-variance divergence with the learnable control variate, and the replay buffer with simulated annealing for off-policy training. We also propose new $SE(3)$ equivariant scale-based parameterization for bias force and relaxed indicator function for frequent training signals. Evaluations on synthetic double-well and real-molecule systems such as Alanine Dipeptide, Polyproline Helix, and Chignolin demonstrate superior accuracy and diversity of our model compared to both classical and ML approaches.
Limitation.
While our algorithm successfully samples transition paths for small proteins, we did not evaluate our algorithm for even larger proteins due to the lack of computational budget. In addition, our work does not generalize across unseen pairs of meta-stable states or different molecular systems. These points to an interesting venue for future research, which would be more appealing for practical applications in drug discovery or material design.
Appendix A Method details
A.1 Log variance formulation
In this section, we derive Equation 5 from Equation 4 to get the explicit expression for log-variance divergence in terms of SDE in Equation 1 and Equation 3. We refer to Nüsken & Richter (2021, Appendix A.1) for the derivation in more general settings.
Our goal is to derive that
$\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}\left[\left(\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}-\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}\left[\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}\right]\right)^{2}\right]=\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}\left[(F_{\bm{v}_{\theta},\tilde{\bm{v}}}-\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}[F_{\bm{v}_{\theta},\tilde{\bm{v}}}])^{2}\right],$ (13)
To this end, we focus on calculating $\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}(\bm{X})$ when $\bm{X}\sim\mathbb{P}_{\tilde{\bm{v}}}$. Following (Nüsken & Richter, 2021, Lemma A.1), we apply Girsanov’s Theorem to calculate the Radon-Nikodym derivative $\frac{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}{\mathrm{d}\mathbb{P}_{0}}$ as follows:
$\frac{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\exp\left(\int^{T}_{0}(\bm{v}_{\theta}^{T}\Sigma^{-1})(\bm{X}_{t})\cdot\mathrm{d}\bm{X}_{t}-\int^{T}_{0}(\Sigma^{-1}\bm{u}\cdot\bm{v}_{\theta})(\bm{X}_{t})\mathrm{d}t-\frac{1}{2}\int^{T}_{0}\lVert\bm{v}_{\theta}(\bm{X}_{t})\rVert^{2}\mathrm{d}t\right)$ (14)
since the state $\bm{X}_{t}$ follows the SDE $\mathrm{d}\bm{X}_{t}=(\bm{u}(\bm{X}_{t})+\Sigma\tilde{\bm{v}}(\bm{X}_{t}))\mathrm{d}t+\Sigma\mathrm{d}\bm{W}_{t}$. We plug it into Equation 14 and utilize the definition of the target path measure $\mathbb{Q}$ in Equation 2 to compute $\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{0}}$ as follows:
$\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}(\bm{X})$ $=\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{0}}\frac{\mathrm{d}\mathbb{P}_{0}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}(\bm{X})$ (15)
$=\log 1_{\mathcal{B}}(\bm{X})-\log Z-\int^{T}_{0}(\bm{v}_{\theta}^{T}\Sigma^{-1})(\bm{X}_{t})\cdot\mathrm{d}\bm{X}_{t}$ (16)
$\quad+\int^{T}_{0}(\Sigma^{-1}\bm{u}\cdot\bm{v}_{\theta})(\bm{X}_{t})\mathrm{d}t+\frac{1}{2}\int^{T}_{0}\lVert\bm{v}_{\theta}(\bm{X}_{t})\rVert^{2}\mathrm{d}t$ (17)
$=\log 1_{\mathcal{B}}(\bm{X})-\log Z-\int^{T}_{0}(\bm{v}_{\theta}\cdot\tilde{\bm{v}})(\bm{X}_{t})\mathrm{d}t$ (18)
$\quad-\int^{T}_{0}\bm{v}_{\theta}(\bm{X}_{t})\cdot\mathrm{d}\bm{W}_{t}+\frac{1}{2}\int^{T}_{0}\lVert\bm{v}_{\theta}(\bm{X}_{t})\rVert^{2}\mathrm{d}t$ (19)
$=F_{\bm{v}_{\theta},\tilde{\bm{v}}}(\bm{X})-\log Z$ (20)
Since $\log Z$ is the constant, it is canceled out in the log-variance divergence as follows:
$\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}\left[\left(\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}-\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}\left[\log\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}\right]\right)^{2}\right]=\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}\left[(F_{\bm{v}_{\theta},\tilde{\bm{v}}}-\mathbb{E}_{\mathbb{P}_{\tilde{\bm{v}}}}[F_{\bm{v}_{\theta},\tilde{\bm{v}}}])^{2}\right],$ (21)
A.2 Connection to existing loss functions on discrete-time domain
In this section, we connect our discretized loss of Equation 8 to the loss function, called relative trajectory balance (Venkatraman et al., 2024, RTB). Like our methods, RTB also amortized inference in target path distribution by training forward distribution on discrete-time domains such as vision, language, and control tasks. When discretized, our loss function is equivalent to the RTB objective.
Our goal is to show that for every paths $\bm{x}_{0:L}$ sampled from the path measure $\mathbb{P}_{\bm{v}_{\bar{\theta}}}$,
$(\hat{F}_{\bm{v}_{\theta},\bm{v}_{\bar{\theta}}}(\bm{x}_{0:L})-w)^{2}=\left(\log\frac{p_{0}(\bm{x}_{0:L})1_{\mathcal{B}}(\bm{x}_{0:L})}{Z_{\theta}p_{\bm{v}_{\theta}}(\bm{x}_{0:L})}\right)^{2},$ (22)
where $w=\log Z_{\theta}$ is a learnable scalar parameter, and path distribution $p_{\bm{v}}(\bm{x}_{0:L})=\prod_{\ell=0}^{L-1}p_{\bm{v}}(\bm{x}_{l+1}|\bm{x}_{l})$ is Markovian, and its transition kernel $p_{\bm{v}}(\bm{x}_{l+1}|\bm{x}_{l})$ are derived from Euler-Maruyama discretization of the SDE in Equation 3 as follows:
$\bm{x}_{l+1}=\bm{x}_{l}+\bm{u}(\bm{x}_{l})\Delta t+\Sigma\bm{v}(\bm{x}_{l})\Delta t+\Sigma\bm{{\epsilon}}_{l},$ (23)
where $\bm{{\epsilon}}_{l}\sim\mathcal{N}(\bm{0},\Delta t)$. To this end, we can calculate as follows:
$\log p_{0}(\bm{x}_{0:L})-\log p_{\bm{v}_{\theta}}(\bm{x}_{0:L})$ (24)
$=\sum_{\ell=0}^{L-1}\log p_{\bm{v}_{\theta}}(\bm{x}_{l+1}|\bm{x}_{l})-\sum_{\ell=0}^{L-1}\log p_{0}(\bm{x}_{l+1}|\bm{x}_{l})$ (25)
$=\frac{1}{2}\sum_{\ell=0}^{L-1}(\Sigma\bm{v}_{\bar{\theta}}\Delta t+\Sigma\bm{{\epsilon}}_{l}-\Sigma\bm{v}_{\theta}\Delta t)^{T}(\Sigma^{T}\Sigma\Delta t)^{-1}(\Sigma\bm{v}_{\bar{\theta}}\Delta t+\Sigma\bm{{\epsilon}}_{l}-\Sigma\bm{v}_{\theta}\Delta t)$ (26)
$-\frac{1}{2}\sum_{\ell=0}^{L-1}(\Sigma\bm{v}_{\bar{\theta}}\Delta t+\Sigma\bm{{\epsilon}}_{l})^{T}(\Sigma^{T}\Sigma\Delta t)^{-1}(\Sigma\bm{v}_{\bar{\theta}}\Delta t+\Sigma\bm{{\epsilon}}_{l})$ (27)
$=\frac{1}{2\Delta t}\sum_{\ell=0}^{L-1}(\lVert\bm{v}_{\bar{\theta}}\Delta t+\bm{{\epsilon}}_{l}-\bm{v}_{\theta}\Delta t\rVert^{2}-\lVert\bm{v}_{\bar{\theta}}\Delta t+\bm{{\epsilon}}_{l}\rVert^{2})$ (28)
$=\frac{1}{2}\sum_{\ell=0}^{L-1}\lVert\bm{v}_{\theta}(\bm{x}_{\ell})\rVert^{2}\Delta t-\sum_{\ell=0}^{L-1}(\bm{v}_{\theta}\cdot\bm{v}_{\bar{\theta}})(\bm{x}_{\ell})\Delta t-\sum_{\ell=0}^{L-1}\bm{v}_{\theta}(\bm{x}_{\ell})\cdot\bm{{\epsilon}}_{\ell}$ (29)
$=\hat{F}_{\bm{v}_{\theta},\bm{v}_{\bar{\theta}}}(\bm{x}_{0:L})-\log 1_{\mathcal{B}}(\bm{x}_{0:L})$ (30)
which implies
$\hat{F}_{\bm{v}_{\theta},\bm{v}_{\bar{\theta}}}(\bm{x}_{0:L})=\log\frac{p_{0}(\bm{x}_{0:L})1_{\mathcal{B}}(\bm{x}_{0:L})}{p_{\bm{v}_{\theta}}(\bm{x}_{0:L})},$ (31)
by subtracting $w$ and squaring both sides, we have
$(\hat{F}_{\bm{v}_{\theta},\bm{v}_{\bar{\theta}}}(\bm{x}_{0:L})-w)^{2}=\left(\log\frac{p_{0}(\bm{x}_{0:L})1_{\mathcal{B}}(\bm{x}_{0:L})}{Z_{\theta}p_{\bm{v}_{\theta}}(\bm{x}_{0:L})}\right)^{2}$ (32)
[…]
A.3 Training algorithm
Algorithm 1 Training
1:  Initialize an empty replay buffer $\hat{\mathcal{D}}$, an policy $\bm{v}_{\theta}$, a scalar parameter $w$, the number of rollout $I$ and training per rollout $J$, and an annealing schedule $\lambda_{\text{start}}=\lambda_{1}>\cdots>\lambda_{I}=\lambda_{\text{end}}$.
2:  for $i=1,\ldots,I$ do
3:     Generate $M$ paths $\{\bm{x}_{0:L}^{(m)}\}^{M}_{m=1}$ from the biased MD simulations with $\bm{v}_{\theta}$ at temperature $\lambda_{i}$.
4:     Update the replay buffer $\hat{\mathcal{D}}\leftarrow\hat{\mathcal{D}}\cup\{\bm{x}_{0:L}^{(m)}\}^{M}_{m=1}$.
5:     for $j=1,\ldots,J$ do
6:        Sample $K$ data $\{\bm{x}_{0:L}^{(k)}\}^{K}_{k=1}$ from $\hat{\mathcal{D}}$.
7:        Update $\theta$ and $w$ with the gradient of $\frac{1}{K}\sum_{k=1}^{K}\left(\log\frac{p_{0}(\bm{x}^{(k)}_{0:L})1_{\mathcal{B}}(\bm{x}^{(k)}_{0:L})}{p_{\bm{v}_{\theta}}(\bm{x}^{(k)}_{0:L})}-w\right)^{2}$
8:     end for
9:  end for
Appendix B Experiment details
B.1 OpenMM configurations
For real-world molecules, Alanine Dipeptide, Polyproline Helix, and Chingolin, we use the VVVR integrator (Sivak et al., 2014) with the step size $\Delta t=1~{}\mathrm{fs}$ and the friction term $\gamma=1~{}\mathrm{ps}^{-1}$. In the TPS-DPS training algorithm, we start simulations at a temperature $\lambda_{\text{start}}=600$K, and end at a temperature $\lambda_{\text{end}}=300$K. Other OpenMM configurations are shown in Table 2.
Table 2: OpenMM configurations for three real-world molecular systems.
Molecule Simulation time ($T$) Force field Solvent
Alanine Dipeptide $1000$ amber99sbildn (Lindorff-Larsen et al., 2010) vaccum
Polyproline Helix $10000$ amber/protein.ff14SBonlysc (Maier et al., 2015) implicit/gbn2
Chignolin $5000$ amber/protein.ff14SBonlysc implicit/gbn2
B.2 Model configurations
We use a 3-layer MLP for the double-well system, and a 6-layer MLP for real-world molecules with ReLU activation functions for neural bias force, potential, and scale. To constrain the output of the neural bias scale parameterization to a positive value, we apply Softplus (Zheng et al., 2015) to the MLP output. As an input to the neural network, we concatenate the current position $(\bm{R}_{t})_{i}$ of the $i$-th atom with its distance to the target position $d_{i}=\lVert(\tilde{\bm{R}}_{\mathcal{B}})_{i}-(\bm{R}_{t})_{i}\rVert_{2}$. For real-world molecules, we apply the Kabsch algorithm (Kabsch, 1976) for heavy atoms to align $\bm{R}_{\mathcal{B}}$ with $\bm{R}_{t}$. We update the parameters of the neural network with a learning rate of $0.0001$, while the scalar parameter $w$ is updated with a learning rate of $0.001$. We clip the gradient norm with 1 to prevent loss from exploding. we train $J=1000$ times per rollout. We report other model configurations in Table 3. For PIPS, we use the model configurations reported by Holdijk et al. (2024). For CVs of SMD, we use backbone dihedral angles $(\phi,\psi)$ for Alanine Dipeptide and RMSD for Polyproline Helix and Chignolin.
Table 3: Model configurations of TPS-DPS.
System # of rollouts ($I$) # of samples ($M$) Batch size ($K$) Buffer size Relaxation ($\sigma$)
Double-well 20 512 512 10000 3
Alanine Dipeptide 1000 16 16 1000 0.1
Polyproline Helix 1000 16 4 200 0.2
Chignolin 1000 16 4 200 0.5
B.3 Evaluation metrics
Root mean square distance (RMSD).
We use the Kabsch algorithm (Kabsch, 1976) for heavy atoms to align the final position with the target position $\bm{R}_{\mathcal{B}}$, using the optimal (proper) rotation and translation to superimpose two heavy atom positions. We calculate RMSD between heavy atoms of the final position and the target position $\bm{R}_{\mathcal{B}}$.
Target hit percentage (THP).
THP measures the success rate of paths arriving at the target meta-stable state $\mathcal{B}$ in a binary manner. Formally, given the final positions $\{R^{(i)}\}^{M}_{i=1}$ of $M$ paths, THP is defined as follows:
$\operatorname{THP}=\frac{|\{i:\bm{R}^{(i)}\in\mathcal{B}\}|}{M}$ (33)
Energy of transition state (ETS).
ETS measures the ability of the method to find probable transition states when crossing the energy barrier. ETS refers to the maximum potential energy among states in a transition path. Formally, given a transition path $\bm{x}_{0:L}$ of length $L$ that reaches the target meta-stable state i.e., $\bm{R}_{L}\in\mathcal{B}$, ETS is defined as follows:
$\operatorname{ETS}(\bm{x}_{0:L})=\max_{\ell\in[0,L]}U(\bm{R}_{\ell})$ (34)
B.4 System details
Double-well potential and dynamics.
Double-well system follows the overdamped Langevin dynamics defined as follows:
$\mathrm{d}\bm{R}_{t}=\frac{-\nabla U(\bm{R}_{t})}{\bm{m}}\mathrm{d}t+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\mathrm{d}\bm{W}_{t}\,.$ (35)
For simplicity, we let $\bm{R}=(x,y)\in\mathbb{R}^{2}$, $\bm{m}=I,\gamma=1,\Delta=0.01,T=10$, and $\lambda=1200$K. To evaluate the ability to find diverse transition paths, we consider the following double-well potential (Hua et al., 2024):
$U(x,y)=\frac{1}{6}\left(4(1-x^{2}-y^{2})^{2}+2(x^{2}-2)^{2}+[(x+y)^{2}-1]^{2}+[(x-y)^{2}-1]^{2}-2\right)\,.$ (36)
This potential has global minima and two saddle points, having two meta-stable states and two reaction channels.
Handedness.
For four points $A,B,C,D$, we define a handedness (Moradi et al., 2009) as follows:
$H_{ABCD}=\frac{\overrightarrow{EF}\cdot(\overrightarrow{CD}\times\overrightarrow{AB})}{|\overrightarrow{EF}|\cdot|\overrightarrow{CD}|\cdot|\overrightarrow{AB}|}.$ (37)
Here, $E$ and $F$ are the midpoints of the vectors $\overrightarrow{AB}$ and $\overrightarrow{CD}$, respectively. With the backbone atoms $X_{1},X_{2},...,X_{N}$ of a Polyproline Helix, we can define the handedness for $N$ atoms as $H=\sum_{i=1}^{N-3}H_{X_{i}X_{i+1}X_{i+2}X_{i+3}}$. In our experiments, we take the alpha carbon $X_{1}$, the carbonyl carbon $X_{2}$ of the first proline residue, the alpha carbon $X_{3}$ of the second residue and alpha carbon $X_{4}$ of third residues to compute the handedness.
Time-lagged independent components (TICA).
[…]
Reproducibility.
We describe experiment details in Appendix B, including detailed simulation configuration and hyper-parameters. In the anonymous link, we provide the code for TPS-DPS.

### v5

@@
 Abstract
-Understanding transition pathways between meta-stable states in molecular systems is crucial to advance material design and drug discovery.
-However, unbiased molecular dynamics simulations are computationally infeasible due to the high energy barriers separating these states.
-Although recent machine learning techniques offer potential solutions, they are often limited to simple systems or rely on collective variables (CVs) derived from costly domain expertise.
-In this paper, we introduce a novel approach that trains diffusion path samplers (DPS) for transition path sampling (TPS) without the need for CVs.
-We recast the problem as an amortized sampling of the target path measure, minimizing the log-variance divergence between the path measure induced by our DPS and the target path measure.
-To ensure scalability for high-dimensional tasks, we introduce (1) a new off-policy training objective based on learning control variates with replay buffers and (2) a scale-based equivariant parameterization of the bias forces.
-We evaluate our approach, coined TPS-DPS, on a synthetic double-well potential and three peptides: Alanine Dipeptide, Polyproline Helix, and Chignolin.
-Results show that our approach produces more realistic and diverse transition pathways compared to existing baselines.
-We also provide links to our project page and code.
+Understanding transition pathways between two meta-stable states of a molecular system is crucial to advance drug discovery and material design.
+However, unbiased molecular dynamics (MD) simulations are computationally infeasible because of the high energy barriers that separate these states.
+Although recent machine learning techniques are proposed to sample rare events, they are often limited to simple systems and rely on collective variables (CVs) derived from costly domain expertise.
+In this paper, we introduce a novel approach that trains diffusion path samplers (DPS) to address the transition path sampling (TPS) problem without requiring CVs.
+We reformulate the problem as an amortized sampling from the transition path distribution by minimizing the log-variance divergence between the path distribution induced by DPS and the transition path distribution.
+Based on the log-variance divergence, we propose learnable control variates to reduce the variance of gradient estimators and the off-policy training objective with replay buffers and simulated annealing techniques to improve sample efficiency and diversity.
+We also propose a scale-based equivariant parameterization of the bias forces to ensure scalability for large systems.
+We extensively evaluate our approach, termed TPS-DPS, on a synthetic system, small peptide, and challenging fast-folding proteins, demonstrating that it produces more realistic and diverse transition pathways than existing baselines.
+We provide links to our project page and code.
+footnotetext: Work done in POSTECH
 1 Introduction
-In material design and drug discovery, it is crucial to understand the mechanisms and kinetics of transitions between meta-stable states of molecular systems, such as protein folding and chemical reactions (Mulholland, 2005; Piana et al., 2012; Ahn et al., 2019; Spotte-Smith et al., 2022).
-Their comprehensive study requires sampling transition paths (Elber, 2016; Lee et al., 2017), which provides insight into mechanisms and energy landscapes.
-However, naïvely sampling transition paths by molecular dynamics (MD) simulations is often computationally intractable due to high energy barriers, which cause an exponential decay in probability to make a transition (Pechukas, 1981).
-To address this problem, researchers have developed enhanced sampling approaches such as steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999), umbrella sampling (Torrie & Valleau, 1977; Kästner, 2011), meta-dynamics (Ensing et al., 2006; Branduardi et al., 2012; Bussi & Branduardi, 2015), on-the-fly probability-enhanced sampling (OPES; Invernizzi & Parrinello, 2020), and adaptive biasing force (ABF; Comer et al., 2015) methods.
-These methods rely on bias forces to facilitate transitions across high energy barriers.
-They are mainly designed based on collective variables (CVs), which are functions of atomic coordinates that capture the slow modes of the transition.
-Although effective for some systems, the reliance on expensive domain knowledge limits the applicability of the methods to systems where CVs are less understood.
-Recently, machine learning has emerged as a promising paradigm for CV-free transition path sampling (Das et al., 2021; Lelièvre et al., 2023; Holdijk et al., 2024).
-The key idea is to parameterize the bias force using a neural network and train the corresponding biased MD simulation to sample realistic transition paths.
+In drug discovery and material design, it is crucial to understand the mechanisms of transitions between meta-stable states of molecular systems, such as protein folding (Salsbury Jr, 2010; Piana et al., 2012), chemical reaction (Mulholland, 2005; Ahn et al., 2019), and nucleation (Thanh et al., 2014; Beaupere et al., 2018).
+Molecular dynamics (MD) simulations have become one of the most widely used tools for sampling these transitions.
+However, sampling transition paths with unbiased MD simulations is computationally expensive due to high energy barriers, which cause an exponential decay in the probability of making a transition (Pechukas, 1981).
+To address this problem, researchers have developed enhanced sampling approaches such as steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999), umbrella sampling (Torrie & Valleau, 1977; Kästner, 2011), meta-dynamics (Ensing et al., 2006; Branduardi et al., 2012; Bussi & Branduardi, 2015), adaptive biasing force (ABF; Comer et al., 2015), on-the-fly probability-enhanced sampling (OPES; Invernizzi & Parrinello, 2020) methods.
+These methods utilize bias forces to facilitate transitions across high energy barriers.
+They require collective variables (CVs), functions of atomic coordinates designed to capture the slow degree of freedom.
+Although effective for simple systems, the reliance on expensive domain knowledge limits the applicability of the methods to complex systems where CVs are less understood.
+Recently, machine learning has emerged as a promising paradigm for CV-free transition path sampling (TPS) (Das et al., 2021; Lelièvre et al., 2023; Holdijk et al., 2024).
+The key idea is to parameterize the bias force using a neural network and train it to sample transition paths directly via the biased MD simulation.
 In particular, Lelièvre et al.
@@
 Das et al.
-(2021) and Holdijk et al.
-(2024) proposed to minimize the Kullback-Leibler (KL) divergence between the path measures induced by the neural network and the target path measure.
-However, minimizing the KL divergence leads to mode collapse, which is hazardous for sampling transition paths due to a lack of exploration.
+(2021); Hua et al.
+(2024); Holdijk et al.
+(2024) considered TPS problem as minimizing the reverse Kullback-Leibler (KL) divergence between the path measures induced by the biased MD and the target path measure.
+However, minimizing the reverse KL divergence suffers from mode collapse, capturing only a subset of modes of the target distribution (Vargas et al., 2023; Richter & Berner, 2024).
 Furthermore, Das et al.
-(2021) and Lelièvre et al.
-(2023) limited their evaluation to low-dimensional synthetic systems.
-Designing machine learning algorithms for CV-free transition path sampling of large molecules remains an open problem.
+(2021); Lelièvre et al.
+(2023); Hua et al.
+(2024); Holdijk et al.
+(2024) limited their evaluation to low-dimensional synthetic systems or small peptides.
+Developing machine learning algorithms that generate accurate and diverse transition pathways for complex molecular systems remains an open challenge.
 Contribution.
-In this work, we propose to solve the transition path sampling problem with the diffusion path sampler (TPS-DPS).
-TPS-DPS trains the bias force to minimize a recently proposed log-variance divergence (Nüsken & Richter, 2021) between the path measure induced by the bias MD and the target path measure.
-To ensure scalability for high-dimensional tasks, e.g., Chignolin folding, we propose a new off-policy training scheme and a scale-based parameterization of the bias forces.
-To be specific, we propose a new training objective to minimize the log-variance divergence, defined between a pair of path measures induced by the biased MD and the unbiased MD conditioned on the target meta-stable state.
-To minimize the variance of gradient estimators and improve sample efficiency, we propose to use (1) learnable control variates and (2) replay buffers with simulated annealing.
-Our choice of minimizing the log-variance is motivated by recent works (Richter et al., 2020; Nüsken & Richter, 2021; Richter & Berner, 2024) that demonstrate many desirable properties of the log-variance divergence such as a robust gradient estimator for training stability and avoidance of differentiation through the SDE solver.
-This allows the use of previously collected samples in a replay buffer, promoting sample efficiency in training and improving the diversity of the generated transition paths by preventing the mode collapse observed in KL divergence.
-We also introduce a new $SE(3)$ equivariant scale-based parameterization for the bias force that guides the biased MD to sample meaningful paths more frequently, based on inductive bias.
-Our key idea is to parameterize the bias force to yield a positive inner product with a direction to decrease the distance between the molecular state and the target meta-stable state, which guarantees the bias force to (locally) decrease the distance.
-We also use the Kabsch algorithm (Kabsch, 1976) to align the molecule and the bias force to the target meta-stable state, guaranteeing $SE(3)$ equivariance for better generalization across molecular states that are identical up to roto-translations.
-We extensively evaluate our method on the synthetic double-well potential and three peptides: Alanine Dipeptide, Polyproline Helix, and Chignolin.
-We compare TPS-DPS with prior ML approach (PIPS; Holdijk et al., 2024), as well as classical non-ML methods, e.g., steered MD (SMD; Izrailev et al., 1999).
+In this work, we propose the diffusion path sampler (DPS) to solve the transition path sampling problem.
+[^1] We coin our method diffusion path sampler since it samples paths using diffusion SDE, similar to diffusion samplers (Zhang & Chen, 2022; Vargas et al., 2023) that use diffusion SDEs for sampling the final state.
+Our approach, coined TPS-DPS, (1) trains the bias force by minimizing a recently proposed log-variance divergence (Nüsken & Richter, 2021) between the path measure induced by the biased MD and the target path measure, and (2) uses scale-based parameterization of the bias force to handle large systems including fast-folding proteins.
+Specifically, to leverage desirable properties of the log-variance divergence, such as robustness of gradient estimator and degree of freedom in reference path measure, we propose to learn a control variate for reducing the variance of gradient estimators and employ off-policy training scheme with replay buffer and simulated annealing to improve sample efficiency and prevent the mode collapse.
+We also introduce a new scale-based equivariant parameterization for the bias force to frequently sample meaningful paths in training.
+Our key idea is to predict the atom-wise positive scaling factor of displacement from current molecular states to the target meta-stable state.
+This guarantees the bias force to decrease the distance between them for every MD step.
+We also use the Kabsch algorithm (Kabsch, 1976) to align the current molecular states with the target meta-stable state, guaranteeing $SE(3)$ equivariance of bias force for better generalization across the states.
+We extensively evaluate our method on the synthetic double-well potential with dual channels, Alanine Dipeptide, and four fast-folding proteins: Chignolin, Trp-cage, BBA, and BBL (Lindorff-Larsen et al., 2011).
+We compare TPS-DPS with the ML approach (PIPS; Holdijk et al., 2024), as well as classical non-ML methods, e.g., steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999).
 Our experiments demonstrate that TPS-DPS consistently generates realistic and diverse transition paths, similar to the ground truth ensemble.
-In addition, ablation studies of algorithmic components verify the effectiveness of our method.
+In addition, we do ablation studies of the proposed components.
 2 Related work
@@
 (1998) proposed shooting methods that use the Markov chain Monte Carlo (MCMC) procedure on path space.
+However, these methods often suffer from long mixing times in path space, which hinders the exploration of diverse transition paths.
+Additionally, high rejection rates can further reduce their efficiency.
 Data driven ML approaches.
-Recently, generative models have been trained to sample a new transition path given a dataset of transition paths.
+Recently, generative models have been used to sample new transition paths given a dataset of transition paths.
 Petersen et al.
@@
 (2023) applied diffusion probabilistic models (Ho et al., 2020) and variational auto-encoders (Kingma & Welling, 2013) for transition path sampling, respectively.
-However, these methods are limited to small systems.
-Concurrent to our work, Jing et al.
-(2024) generated molecular trajectories by task-wise conditions given an initial state and viewed transition path sampling as an interpolation task between two molecular states.
-We also note some related works that are not directly comparable to our method.
+However, these methods are limited to small systems and struggle to collect data using unbiased MD simulations due to the high energy barriers.
 Klein et al.
-(2024) proposed to accelerate MD simulation through coarse-grained path generation that may include multiple meta-stable states.
-However, this work is incapable of targeting a specific meta-stable state and generating fine-grained transition paths.
+(2024); Schreiner et al.
+(2024); Jing et al.
+(2024) proposed to accelerate MD using time-coarsened dynamics, but the time-coarsened dynamics cannot capture the details of the transition, e.g., the transition states.
 Duan et al.
 (2023); Kim et al.
-(2024) use neural networks to generate transition states of a given chemical reaction, but cannot generate transition paths.
+(2024) used neural networks to generate transition states of a given chemical reaction, but cannot generate transition paths.
 Data free ML approaches.
@@
 (2021); Lelièvre et al.
+(2023); Sipka et al.
 (2023); Hua et al.
 (2024); Holdijk et al.
-(2024) trained generators to sample transition paths using biased MD.
+(2024) trained the bias forces to sample transition paths using the biased MD.
 Lelièvre et al.
-(2023) used reinforcement learning to train the generator but focused on escaping an initial meta-stable state rather than reaching the target meta-stable state.
-Next, Holdijk et al.
-(2024) considered transition path sampling as minimizing the reverse KL divergence between path distribution from biased MD and transition path distribution.
-In this work, we mainly compare our method with Holdijk et al.
-(2024) and provide quantitative analysis for real-world molecule systems.
-We also note that, concurrent to our work, Du et al.
-(2024) also considered variational formulation of the TPS problem using Doob’s $h$-transform.
+(2023) used reinforcement learning to train the bias forces but focused on escaping an initial meta-stable state rather than targeting a given meta-stable state.
+Sipka et al.
+(2023) used differentiable biased MD simulation to train bias potential and introduce partial back-propagation and graph mini-batching techniques to resolve computational issues in differentiable simulation.
+Das et al.
+(2021); Hua et al.
+(2024); Holdijk et al.
+(2024) considered the TPS problem as minimizing the reverse KL divergence between path distribution from biased MD and transition path distribution.
+Das et al.
+(2021); Hua et al.
+(2024) limited their evaluation to low-dimensional synthetic systems.
+In this work, we mainly compare our method with (PIPS; Holdijk et al., 2024).
+Concurrent to our work, Du et al.
+(2024) considered the TPS problem as minimizing Doob’s Lagrangian objective with boundary constraints.
+They parameterized marginal distribution as (mixture) Gaussian path distribution to satisfy the boundary constraints without simulation in training time and sampled transition paths with the bias force derived from the Fokker-Planck equation in inference time.
 3 Transition path sampling with diffusion path samplers
+In this section, we introduce our method, coined transition path sampling with diffusion path sampler (TPS-DPS).
+Our main idea is to formulate the transition path sampling (TPS) problem as a minimization of the log-variance divergence (Nüsken & Richter, 2021) between two path measures: the path measure induced by DPS and that of transition paths.
+Our main methodological contribution is twofold: (1) a new off-policy training algorithm that minimizes the log-variance divergence with the learnable control variate, replay buffer, and simulated annealing (2) a $SE(3)$ equivariant scale-based parameterization of the bias force that has an inductive bias for dense training signals in large systems.
 Figure 1: Problem setup.
-Potential energy landscape of Alanine Dipeptide, where the sampled path from state $\mathcal{A}$ to state $\mathcal{B}$ is highlighted in yellow dotted lines.
-We visualize snapshots in the transition path highlighted in white circles and the transition state highlighted in the white star.
-In this section, we introduce our method, coined transition path sampling with diffusion path sampling (TPS-DPS).
-Our main idea is to formulate the transition path sampling (TPS) problem as a minimization of log-variance divergence (Nüsken & Richter, 2021) between two path measures: the path measure of transition paths and that of biased MD simulations parameterized by a neural network.
-Our main methodological contribution is twofold: (1) a new training algorithm that minimizes the log-variance divergence with a learned control variate and replay buffer, and (2) a $SE(3)$ equivariant scale-based parameterization of the biased MD that provides inductive bias for dense training signals in high-dimensional problems.
+The sampled transition path (yellow dotted lines) from the state $\mathcal{A}$ to the state $\mathcal{B}$ on the free energy landscape of Alanine Dipeptide.
+We visualize the snapshots (white circles) of the transition path and the transition state (white star).
 3.1 Problem setup
-Our goal is to sample transition paths from one meta-stable state to another meta-stable state given a molecule system.
+Our goal is to sample transition paths from one meta-stable state to another meta-stable state given a molecular system.
 We provide an example of the problem for Alanine Dipeptide in Figure 1.
-We view this as a task to sample paths from an unbiased molecular dynamics (MD), described by a stochastic differential equation (SDE), conditioned on its starting and ending points.
+We view this as a task to sample paths from an unbiased molecular dynamics (MD) in Equation 1 conditioned on its starting and ending points of initial and target meta-stable states, respectively.
 To solve this task, we train the bias force parameterized by a neural network to amortize the sampling procedure.
 Molecular dynamics.
-We consider a MD simulation on time interval $[0,T]$ that describes motion of a molecular state $\bm{X}_{t}=(\bm{R}_{t},\bm{V}_{t})\in\mathbb{R}^{6N}$ at time $t$ where $N$ is the number of atoms, $\bm{R}_{t}\in\mathbb{R}^{3N}$ is the atom-wise positions and $\bm{V}_{t}\in\mathbb{R}^{3N}$ is the atom-wise velocities.
+We consider a MD simulation on time interval $[0,T]$, i.e., the motion of a molecular state $\bm{X}_{t}=(\bm{R}_{t},\bm{V}_{t})\in\mathbb{R}^{6N}$ at time $t$ where $N$ is the number of atoms, $\bm{R}_{t}\in\mathbb{R}^{3N}$ is the atom-wise positions and $\bm{V}_{t}\in\mathbb{R}^{3N}$ is the atom-wise velocities.
 In particular, we adopt Langevin dynamics (Bussi & Parrinello, 2007) defined as the following SDE:
@@
 where $U$, $\bm{m}$, $\gamma$, $k_{B}$, $\lambda$, and $\bm{W}_{t}$ denote the potential energy function, the atom-wise masses, the friction term, the Boltzmann constant, the absolute temperature, and the Brownian motion, respectively, and $\bm{\zeta}\in\mathbb{R}^{3N}$ is a vector of positive infinitesimal values.
-We let $\mathbb{P}_{0}$ denote the path measure induced by Equation 1 in the space $\mathcal{C}([0,T];\mathbb{R}^{6N})$ of continuous functions $\bm{X}:[0,T]\rightarrow\mathbb{R}^{6N}$.
+MD in Equation 1 induces the path measure, denoted by $\mathbb{P}_{0}$, which refers to the positive measure defined on measurable subsets of the path space $\mathcal{C}([0,T];\mathbb{R}^{6N})$ consisting of continuous functions $\bm{X}:[0,T]\rightarrow\mathbb{R}^{6N}$.
+The path (probability) measure $\mathbb{P}_{0}$ induced by MD assigns high probability to a set of the probable paths when solving MD.
 Transition path sampling.
@@
 Since the meta-stable state $\mathcal{A}$ and $\mathcal{B}$ are not well-specified for many molecular systems, we simplify this task by (1) fixing a local minima $\bm{R}_{\mathcal{A}},\bm{R}_{\mathcal{B}}$ of the potential energy function in the meta-stable states $\mathcal{A},\mathcal{B}$ and (2) sampling a transition path $\bm{X}$ that starts from the state $\bm{R}_{0}=\bm{R}_{\mathcal{A}}$ and ends at the vicinity of $\bm{R}_{\mathcal{B}}$.
-To be specific, we aim to sample from the target path measure $\mathbb{Q}$ which reweights the path measure $\mathbb{P}_{0}$ via the Radon–Nikodym derivative, defined as follows:
-$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\frac{1_{\mathcal{B}}(\bm{X})}{\mathbb{E}_{\mathbb{P}_{0}}\left[1_{\mathcal{B}}(\bm{X})\right]},\quad 1_{\mathcal{B}}(\bm{X})=\begin{cases}1&\text{if}~{}\lVert\bm{R}_{T}-\rho_{T}\cdot\bm{R}_{\mathcal{B}}\rVert\leq\delta,\\ 0&\text{otherwise},\end{cases}$ (2)
-where $\cdot$ denotes group action associated with the $SE(3)$ space and $\rho_{T}\cdot\bm{R}_{\mathcal{B}}$ is the aligned position by the optimal roto-translation $\rho_{T}\in SE(3)$ to minimize its Euclidean distance to $\bm{R}_{T}$, i.e., $\rho_{T}=\operatorname*{arg\,min}_{\rho\in SE(3)}\lVert\bm{R}_{T}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$.
+To be specific, we aim to sample from the target path measure $\mathbb{Q}$, which is obtained by reweighting the path measure $\mathbb{P}_{0}$ with the (normalized) indicator function.
+The indicator function assigns zero weight to paths that do not reach the vicinity of the target position $\bm{R}_{\mathcal{B}}$.
+Formally, the reweighting function is called the Radon–Nikodym derivative defined as follows:
+$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\frac{1_{\mathcal{B}}(\bm{X})}{Z},\quad 1_{\mathcal{B}}(\bm{X})=\begin{cases}1&\text{if}~{}\lVert\rho_{T}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{T}\rVert\leq\delta,\\ 0&\text{otherwise},\end{cases}\quad Z=\mathbb{E}_{\mathbb{P}_{0}}\left[1_{\mathcal{B}}(\bm{X})\right],$ (2)
+where $\cdot$ denotes group action associated with the $SE(3)$ space and $\rho_{T}\cdot\bm{R}_{\mathcal{B}}$ is the aligned target position by the optimal roto-translation $\rho_{T}\in SE(3)$ to minimize its Euclidean distance to $\bm{R}_{T}$, i.e., $\rho_{T}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\bm{R}_{T}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$.
 Such a transformation can be obtained from the Kabsch algorithm in $O(N)$ complexity (Kabsch, 1976).
 Note that one may consider naïve rejection sampling to sample transition paths, based on running unbiased MD to sample a path $\bm{X}$ from the path measure $\mathbb{P}_{0}$ and accepting if the path $\bm{X}$ arrives at the neighborhood of the position $\bm{R}_{\mathcal{B}}$ with the radius $\delta$.
-However, this method does not scale to high-dimensional or low-temperature problems, since the sampled path by unbiased MD rarely reaches the target states due to the high energy barriers, i.e., the rejection ratio is too high.
+However, this method struggles to sample transition paths of systems with high energy barriers since the sampled path by unbiased MD rarely reaches the target states, i.e., the rejection ratio is too high.
 3.2 Log-variance minimization
 In this section, we propose our algorithm to amortize transition path sampling.
-Our key idea is to train a neural network to induce a path measure that matches the target path measure $\mathbb{Q}$, using the log-variance divergence (Nüsken & Richter, 2021) between the path measures.
+Our key idea is to train a neural network to induce a path measure that matches the target path measure $\mathbb{Q}$, using the log-variance divergence between the path measures.
 We propose a new training scheme to minimize the log-variance divergence based on learning the control variate of its gradient and a replay buffer to improve sample efficiency and diversity.
@@
 We provide the full derivation in Section A.1.
-Compared to KL divergence, the log-variance divergence provides a robust gradient estimator and avoids differentiating through the SDE solver.
-(Richter et al., 2020; Nüsken & Richter, 2021).
+Compared to KL divergence, the log-variance divergence provides a robust gradient estimator and avoids differentiating through the SDE solver (Richter et al., 2020; Nüsken & Richter, 2021).
 Minimizing with learnable control variate.
-To minimize the log-variance divergence, we consider the following loss function that replaces the estimation of $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ by learning a parameter $w$:
+To minimize the log-variance divergence, we consider the following loss that replaces the estimation of $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ by learning a scalar parameter $w$:
 $\mathcal{L}(\theta,w)=\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}\left[(F_{\bm{v}_{\theta},\bm{v}_{\theta}}-w)^{2}\right],$ (7)
@@
 When optimized, the control variate $w$ estimates the expectation $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ since $\operatornamewithlimits{argmin}_{w}\mathcal{L}(\theta,w)=\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$.
-Thus, jointly optimizing $(\theta,w)$ with the gradient step can be interpreted as jointly minimizing log-variance divergence and estimating $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ using $w$.
-Off-policy training with replay buffer.
+Thus, jointly optimizing $(\theta,w)$ with the gradient step can be interpreted as jointly minimizing log-variance divergence and estimating $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ utilizing $w$.
+Off-policy training with replay buffer and simulated annealing.
 To leverage the degree of freedom in reference path measure for the log-variance divergence, we allow discrepancy between reference path measure and current path measure, called off-policy training, which is widely used in discrete-time reinforcement learning (Mnih et al., 2013; Bengio et al., 2021).
 For the sample efficiency, we reuse the samples with a replay buffer $\mathcal{D}$ which stores path samples from the path measure $\mathbb{P}_{\bm{v}_{\bar{\theta}}}$ associated with previous policies $\bm{v}_{\bar{\theta}}$.
-Our modified loss function $\mathcal{L}^{\mathcal{D}}$ with replay buffer $\mathcal{D}$ is defined as follows:
+Our modified loss function $\mathcal{L}^{\mathcal{D}}$ with $\mathcal{D}$ is defined as follows:
 $\mathcal{L}^{\mathcal{D}}(\theta,w)=\mathbb{E}_{(\bm{v}_{\bar{\theta}},\bm{X})\sim\mathcal{D}}[(F_{\bm{v}_{\theta},\bm{v}_{\bar{\theta}}}(\bm{X})-w)^{2}].$ (8)
-Using the replay buffer also prevents mode collapse, using diverse paths from different path measures.
-Similar to other off-policy training algorithms (Malkin et al., 2022; Kim et al., 2023), we use simulated annealing to collect diverse paths that cross high energy barriers.
+The replay buffer also prevents mode collapse, using diverse paths from different path measures.
+Additionally, in line with other off-policy training algorithms (Malkin et al., 2022; Kim et al., 2023), we utilize the simulated annealing technique to sample diverse paths that cross high-energy barriers.
 Discretization.
@@
 (2024).
-We describe our training algorithm in Section A.3.
+Algorithm 1 Training
+1:  Initialize an empty replay buffer $\hat{\mathcal{D}}$, an policy $\bm{v}_{\theta}$, a scalar parameter $w$, the number of rollout $I$ and training per rollout $J$, and an annealing schedule $\lambda_{\text{start}}=\lambda_{1}>\cdots>\lambda_{I}=\lambda_{\text{end}}$.
+2:  for $i=1,\ldots,I$ do
+3:     Generate $M$ paths $\{\bm{x}_{0:L}^{(m)}\}^{M}_{m=1}$ from the biased MD simulations with $\bm{v}_{\theta}$ at temperature $\lambda_{i}$.
+4:     Update the replay buffer $\hat{\mathcal{D}}\leftarrow\hat{\mathcal{D}}\cup\{\bm{x}_{0:L}^{(m)}\}^{M}_{m=1}$.
+5:     for $j=1,\ldots,J$ do
+6:        Sample $K$ data $\{\bm{x}_{0:L}^{(k)}\}^{K}_{k=1}$ from $\hat{\mathcal{D}}$.
+7:        Update $\theta$ and $w$ with the gradient of $\frac{1}{K}\sum_{k=1}^{K}\left(\log\frac{p_{0}(\bm{x}^{(k)}_{0:L})1_{\mathcal{B}}(\bm{x}^{(k)}_{0:L})}{p_{\bm{v}_{\theta}}(\bm{x}^{(k)}_{0:L})}-w\right)^{2}$.
+8:     end for
+9:  end for
+We describe our training algorithm in Algorithm 1.
 Overall, our off-policy training algorithm iterates through four steps: (1) sampling paths from the biased MD simulation with current policy $\bm{v}_{\theta}$ at high temperature, (2) storing sampled paths in the replay buffer $\hat{\mathcal{D}}$, (3) sampling a batch of the paths from the replay buffer, and (4) training current policy $\bm{v}_{\theta}$ by minimizing the loss in Equation 10.
-After minimization, biased MD simulation can directly sample transition paths by amortizing inference in the target path measure.
-3.3 Parameterization for high dimensional tasks
-In this section, we introduce new parameterizations of the bias force and the indicator function for high-dimensional tasks.
-Our parameterization is designed around alleviating the problem of sparse training signal, where the model is unable to learn anything due to the failure of collecting meaningful paths that end near the meta-stable state.
-This problem is especially severe in high-dimensional problems such as large molecules.
+After training, the biased MD simulation can directly sample transition paths from the target path measure.
+3.3 Parameterization for large systems
+In this section, we introduce new parameterizations of the bias force and the indicator function to sample transition paths of large systems.
+Our parameterization is designed to alleviate the problem of sparse training signal, where the model struggles to collect meaningful paths that end at the vicinity of the target meta-stable state in training.
+This problem is especially severe in large systems.
 Bias force parameterization.
-To frequently sample the meaningful paths, we parameterize the bias force to guarantee (locally) decreasing the distance between the current molecular state and the target meta-stable state.
-This is done by parameterizing the bias force as positive atom-wise scaling of the force minimizing the distance, which implies that the bias force can be decomposed into two terms: (1) the direction to decrease the Euclidean distance and (2) the direction that leaves the Euclidean distance invariant.
-The bias force also satisfies roto-translational equivariance to the current state $\bm{X}_{t}$, aligning with the symmetry of the transition path sampling problem for better generalization.
-To be specific, we use a roto-translation $\rho_{t}\in SE(3)$ to align $\bm{R}_{\mathcal{B}}$ with $\bm{R}_{t}$, as we do for indicator function in Equation 2.
-To achieve $SE(3)$ equivariant, we parameterize the bias force as follows:
+To frequently sample the meaningful paths, we aim to parameterize the bias force which guarantees to reduce the distance between the current molecular state and the target meta-stable state for every MD step.
+This is achieved by predicting the atom-wise positive scaling factor of the direction to the aligned target meta-stable state rather than predicting force or potential directly.
+Moreover, we design the bias force to satisfy roto-translational equivariance to the current molecular state input $\bm{X}_{t}$, aligning with the symmetry of the transition path sampling problem for better generalization and parameter efficiency.
+To be specific, we achieve the inductive bias for dense training signals and the $SE(3)$ equivariance to the current molecular state input $\bm{X}_{t}$ by parameterizing the bias force as follows:
 $\bm{b}(\bm{X}_{t})=\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))(\rho_{t}\cdot{\bm{R}}_{\mathcal{B}}-\bm{R}_{t}),$ (11)
-where $\bm{s}_{\theta}(\cdot)\in\mathbb{R}_{+}^{3N}$ is a neural network constrained to have positive output elements.
-One can observe the positive inner product between the bias force and the direction to minimize the Euclidean distance, i.e., $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$, as follows:
-$\bm{b}(\bm{X}_{t})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})=(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})^{T}\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})\geq 0,$ (12)
-which implies that the bias force can be decomposed into a vector positively aligning with $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$ and a vector orthogonal to $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$.
-We also note that the bias force is $SE(3)$ equivariant to the current state $\bm{X}_{t}$ since it is the product of an invariant term $\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))$ and an equivariant term $(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$.
-This stems from the fact that the Kabsch algorithm is equivariant to the current state $\bm{X}_{t}$, i.e., roto-translating position $\bm{R}_{t}\mapsto\bar{\rho}\cdot\bm{R}_{t}$ leads to roto-translating output of the Kabsch algorithm $\rho_{t}\mapsto\bar{\rho}\cdot\rho_{t}$.
-In the experiments, we also consider other equivariant parameterizations that are less constrained: (1) directly predicting the bias force by $\rho_{t}\cdot\bm{b}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}^{3N}$ and (2) predicting the force as gradient of the invariant bias potential $\nabla b_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}$.
-We observe the parameterizations to be useful for low-dimensional tasks, but suffer from the sparse training signal on high-dimensional tasks.
+where $\bm{s}_{\theta}(\cdot)\in\mathbb{R}_{+}^{3N}$ is a neural network constrained to have positive output elements and predicts atom-wise scaling factors and the optimal roto-translation $\rho_{t}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\bm{R}_{t}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$ which aligns $\bm{R}_{\mathcal{B}}$ with $\bm{R}_{t}$, as in Equation 2.
+We note that the bias force (divided by atom-wise masses) is positively correlated with the direction to target state, i.e., $(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})>0$.
+To formalize the benefit of the positive correlation between the bias force and the direction to the target state, one can prove that there always exists a small enough step size $\Delta t$ that decreases the distance between the current state $\bm{R}_{t}$ and the aligned target state $\rho^{\prime}_{t+\Delta t}\cdot{\bm{R}}_{\mathcal{B}}$, i.e.,
+$\lVert\rho^{\prime}_{t+\Delta t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t+\Delta t}^{\prime}\rVert<\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert,$ (12)
+where $\bm{R}_{t+\Delta t}^{\prime}=\bm{R}_{t}+\bm{b}(\bm{X}_{t})\Delta t/\bm{m}$ is the position updated by the bias force with step size $\Delta t$ and $\rho^{\prime}_{t+\Delta t}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\bm{R}^{\prime}_{t+\Delta t}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$.
+We provide the proof of Equation 12 in Section A.3.
+In the experiments, we also consider other equivariant parameterizations that are less constrained: (1) directly predicting the equivariant bias force $\bm{b}(\bm{X}_{t})=\rho_{t}\cdot\bm{b}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}^{3N}$ and (2) predicting the invariant bias potential $b_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}$ to obtain the bias force $\bm{b}(\bm{X}_{t})=-\nabla b_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})\in\mathbb{R}^{3N}$.
+We observe these two parameterizations to be useful for low-dimensional tasks but struggle to produce meaningful paths in large systems during training.
+As shown in Figure 2, bias forces with the positive scaling parameterization are positively correlated with the direction to the target position regardless of network parameters, unlike direct force parameterizations.
+(a) Direct prediction
+(b) Positive scaling
+Figure 2: Visualization of the bias force fields of two different bias force parameterizations with initialized neural networks.
+(2(a)) directly predicting the bias force and (2(b)) predicting the positive scaling factors of direction to the target position (white circle).
 Indicator function parameterization.
-We also propose to relax the indicator function $1_{\mathcal{B}}$ as a radial basis function (RBF) kernel $\tilde{1}_{\mathcal{B}}(\bm{X})=k(\bm{R}_{T},\rho^{-1}\cdot{\bm{R}}_{\mathcal{B}};\sigma^{2})$ which measures the similarity between two positions and $\sigma>0$ controls the degree of relaxation.
+We propose to relax the indicator function $1_{\mathcal{B}}$ into a radial basis function (RBF) kernel $\tilde{1}_{\mathcal{B}}(\bm{X})=k(\bm{R}_{T},\rho^{-1}\cdot{\bm{R}}_{\mathcal{B}};\sigma^{2})$ which measures the similarity between two positions where $\sigma>0$ controls the degree of relaxation.
 The range of RBF kernel $k$ is bounded by the interval $(0,1]$ so that $\log\tilde{1}_{\mathcal{B}}(\bm{X})$ is well-defined and $\tilde{1}_{\mathcal{B}}(\bm{X})$ represents the binary indicator function smoothly.
-To capture high training signal, we focus on the signal from subtrajectories of sampled paths, taking maximum over RBF kernel values of all intermediate states by $\tilde{1}_{\mathcal{B}}^{\text{max}}(\bm{X})=\max_{t\in[0,T]}k(\bm{R}_{t},\tilde{\bm{R}}_{\mathcal{B}};\sigma^{2})$.
+To capture a high training signal from subtrajectories of sampled paths, we propose to take maximum over RBF kernel values of all intermediate states by $\tilde{1}_{\mathcal{B}}^{\text{max}}(\bm{X})=\max_{t\in[0,T]}k(\bm{R}_{t},\tilde{\bm{R}}_{\mathcal{B}};\sigma^{2})$.
 To extract the subtrajectory with a high training signal, we can truncate the paths at the time that maximizes RBF kernel values, allowing variable path lengths.
-Also, the relaxed indicator function is $SE(3)$ invariant to $\bm{R}_{t}$ due to the Kabsch algorithm.
+Notably, the relaxed indicator function is $SE(3)$ invariant to $\bm{R}_{t}$ because of the Kabsch algorithm.
 4 Experiment
-Table 1: Benchmark scores on the double-well system and three real-world molecular systems: Alanine Dipeptide, Polyproline Helix, and Chignolin.
-All metrics are averaged over 1024 paths for the double-well system, and 64 paths for real-world molecular systems.
+Table 1: Benchmark scores on the double-well system and Alanine Dipeptide.
+All metrics are averaged over 1024 paths for the double-well system, and 64 paths for Alanine Dipeptide.
 ETS is computed for paths that hit the target meta-stable state, and the best results are highlighted in bold.
-TPS-DPS predicting the bias force, potential, and atom-wise scaling are denoted by (F), (P), and (S), respectively.
+Predicting the bias force, potential, and atom-wise positive scaling are denoted by (F), (P), and (S), respectively.
 UMD ($\lambda$) denotes unbiased MD with temperature $\lambda$ and SMD ($k$) denotes steered MD with the force constant $k$.
-Unless otherwise specified, paths are generated by MD simulation at 1200K for double-well and 300K for real-world molecules.
+Unless otherwise specified, paths are generated by MD simulation at 1200K for double-well and 300K for Alanine Dipeptide.
 Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
@@
 TPS-DPS (S, Ours) 0.01 $\pm$ 0.03 99.80 1.73 $\pm$ 0.20
-Polyproline Helix
-UMD (300K) 2.22 $\pm$ 0.11 0.00 -
-UMD (1200K) 1.38 $\pm$ 0.45 10.94 1010.28 $\pm$ 38.44
-SMD (5k) 1.68 $\pm$ 0.17 54.69 350.58 $\pm$ 14.36
-SMD (10k) 1.26 $\pm$ 0.06 100.00 355.62 $\pm$ 14.83
-PIPS (F) 2.64 $\pm$ 0.15 0.00 -
-PIPS (P) 1.85 $\pm$ 0.13 93.75 574.66 $\pm$ 20.49
-TPS-DPS (F, Ours) 1.53 $\pm$ 0.12 98.44 418.17 $\pm$ 45.54
-TPS-DPS (P, Ours) 1.35 $\pm$ 0.12 100.00 345.00 $\pm$ 32.58
-TPS-DPS (S, Ours) 1.17 $\pm$ 0.02 100.00 342.00 $\pm$ 20.28
 Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
@@
 Alanine Dipeptide
-UMD (300K) 1.59 $\pm$ 0.15 0.00 -
-UMD (3600K) 1.19 $\pm$ 0.32 6.25 812.47 $\pm$ 148.80
-SMD (10) 0.86 $\pm$ 0.21 7.81 33.15 $\pm$ 6.46
+UMD (300K) 1.59 $\pm$ 0.15 $0$
+UMD (3600K) 1.19 $\pm$ 0.32 06.25 812.47 $\pm$ 148.80
+SMD (10) 0.86 $\pm$ 0.21 07.81 33.15 $\pm$ 6.460
 SMD (20) 0.56 $\pm$ 0.27 54.69 78.40 $\pm$ 12.76
 PIPS (F) 0.66 $\pm$ 0.15 43.75 28.17 $\pm$ 10.86
-PIPS (P) 1.66 $\pm$ 0.03 0.00 -
+PIPS (P) 1.66 $\pm$ 0.03 $0$
 TPS-DPS (F, Ours) 0.16 $\pm$ 0.06 92.19 19.82 $\pm$ 15.88
@@
 TPS-DPS (S, Ours) 0.25 $\pm$ 0.20 76.00 22.79 $\pm$ 13.57
+In this section, we compare our method, termed TPS-DPS, with both classical non-ML and ML approaches, assessing the accuracy and diversity of sampled transition paths.
+We begin with the double-well system and Alanine Dipeptide followed by the four fast-folding proteins: Chignolin, Trp-cage, BBA, and BBL.
+Additionally, we conduct ablation studies to validate the effectiveness of each component in our method.
+All real-world molecular systems are simulated using the OpenMM library (Eastman et al., 2023).
+Details on OpenMM simulation and model configurations are provided in Sections B.1 and B.2, respectively.
+In Appendix C, we analyze the time complexity of TPS-DPS and evaluate the number of energy evaluations and runtime in training and inference time.
+Evaluation Metrics.
+We consider three metrics to evaluate models: (Kabsch) RMSD, THP, and ETS.
+The root mean square distance (RMSD) measures the ability of the model to produce final positions of paths close to the target position $\bm{R}_{\mathcal{B}}$.
+The target hit percentage (THP) measures the ability of the model to produce final positions of paths that successfully arrive at the target meta-stable state $\mathcal{B}$.
+Finally, the energy of the transition state (ETS) measures the ability of the model to identify probable transition states.
+For further details, refer to Section B.3.
+Baselines.
+We compare TPS-DPS with both non-ML and ML baselines.
+For non-ML baselines, we consider unbiased MD (UMD) with various temperatures and steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999) with various force constants and collective variables (CVs).
+For ML baselines, we consider a CV-free transition path sampling method, path integral path sampling (PIPS; Holdijk et al., 2024) which also trains a bias force by minimizing the KL divergence between path measures induced by the biased MD and the target path measure.
+(a) Potential energy landscapes
+(b) Distributions of the energy and $y$
+Figure 3: Visualization of potential energy landscapes and distributions in double-well potential.
+(3(a)) Visualization of the learned bias potential $b_{\theta}$ of TPS-DPS (P).
+(3(b)) Distributions of the potential energy and $y$ coordinate of transition states from 1024 transition paths sampled by each method.
+4.1 Double-well system
+UMD
+SMD ($0.5$)
+SMD ($1$)
+TPS-DPS (F)
+TPS-DPS (P)
+Ground Truth
+UMD
+SMD (10)
+SMD (20)
+PIPS (F)
+TPS-DPS (F)
+TPS-DPS (P)
+Figure 4: Visualization of sampled paths on energy landscapes.
+For the double-well system, we aim to sample transition paths from the left meta-stable state to the right on the potential energy landscape (top).
+For Alanine Dipeptide, we aim to sample conformational changes from the $C5$ (upper left) to the $C7ax$ (lower right) on the Ramachandran plot (bottom).
+White circles and stars indicate meta-stable states and saddle points, respectively.
+We begin by evaluating our method on a synthetic two-dimensional double-well system with dual channels.
+This system has two global minima representing the meta-stable states and two reaction pathways with the corresponding saddle points.
+We aim to sample transition paths from the left state $\bm{R}_{\mathcal{A}}$ to the right meta-stable states =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\bm{R}-\bm{R}_{\mathcal{B}}\rVert<0.5\}$.
+We collect ground truth path ensembles by rejection sampling which proposes paths sampled from the unbiased MD simulations and accepts if the final states are in the target meta-stable states $\mathcal{B}$.
+We provide more details on the double-well system in Section B.4.
+As shown in Table 1, Figure 3, and Figure 4, TPS-DPS outperforms baselines regardless of the bias force parameterizations and generates more similar transition paths to the ground truth than baselines.
+In Figure 3, the bias potential accelerates the transition by increasing the potential energy near the initial meta-stable state while decreasing the potential energy near the two energy barriers.
+Moreover, the distribution of energy and $y$ coordinate of the transition states from TPS-DPS is closest to the ground truth compared with other baselines, successfully capturing two reaction channels.
+In Figure 4, TPS-DPS (F) and (P) generate similar transition paths to the ground truth while UMD at 1200K fails to escape the initial state and SMD struggles to pass the saddle points.
+4.2 Alanine Dipeptide
+For real-world molecules, we first consider Alanine Dipeptide consisting of two alanine residues and aim to sample conformational changes from the meta-stable state $C5$ to $C7ax$ as shown in Figure 4.
+The target meta-stable states are defined as =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$, where $\xi(\bm{R})=(\phi,\psi)$ is a well-known collective variable which consists of two backbone dihedral angles.
+Alanine Dipeptide has two reaction channels with the corresponding saddle points.
+In Table 1 and Figure 4, our method shows superior performance regardless of the bias force parameterizations and successfully generates diverse transition paths that capture two reaction channels.
+Compared to our method, UMD at 300K fails to escape the initial state, SMD with the two backbone torsion CV generates transition paths with less probable transition states, and two-way shooting struggles to find plausible transition states.
+PIPS generates transition paths with only one reaction channel, suffering from mode collapse.
+Table 2: Benchmark scores on fast-folding proteins. All metrics are averaged over 64 paths. Unless otherwise specified, paths are generated at 300K for Chignolin and 400K for the others.
+Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
+Å % $\mathrm{kJmol}^{-1}$
 Chignolin
@@
 UMD (1200K) 7.23 $\pm$ 0.93 1.56 388.17
-SMD (10k) 1.26 $\pm$ 0.31 6.25 -527.95 $\pm$ 93.58
-SMD (15k) 1.17 $\pm$ 0.31 23.44 -237.15 $\pm$ 122.29
+SMD (10k) 1.26 $\pm$ 0.31 6.25 -527.95 $\pm$ 93.580
+SMD (20k) 0.85 $\pm$ 0.24 34.38 179.52 $\pm$ 138.87
 PIPS (F) 4.66 $\pm$ 0.17 0.00 -
@@
 TPS-DPS (P, Ours) 3.87 $\pm$ 0.42 0.00 -
-TPS-DPS (S, Ours) 1.17 $\pm$ 0.66 59.38 -780.18 $\pm$ 216.93
-In this section, we compare our method, called TPS-DPS, with both classical and ML approaches, assessing the accuracy and diversity of sampled transition paths.
-We begin with a synthetic double-well system at 1200K, followed by three real-world molecular systems with various numbers of amino acids: Alanine Dipeptide, Polyproline Helix, and Chignolin.
-Additionally, we conduct ablation studies to validate the effectiveness of each component in our method.
-All real-world molecular systems are simulated using the OpenMM library (Eastman et al., 2023).
-Details on OpenMM simulation and model configurations are provided in Sections B.1 and B.2, respectively.
-Evaluation Metrics.
-We consider three metrics for evaluation: RMSD, THP, and ETS.
-The root mean square distance (RMSD) calculates the Euclidean distance between the atoms in the final position of the sampled path and the target position $\bm{R}_{\mathcal{B}}$, with the final position aligned to the target.
-The target hit percentage (THP) measures the number of paths that successfully arrive at the target meta-stable state $\mathcal{B}$.
-Finally, the energy of the transition state (ETS) measures the ability to identify probable transition states.
-For further details, refer to Section B.3.
-Baselines.
-We compare TPS-DPS with both non-ML and ML baselines.
-For non-ML baselines, we consider unbiased MD (UMD) in various temperatures and steered MD (SMD; Izrailev et al., 1999) with various force constants $k$ and collective variables.
-For ML baselines, we consider a CV-free transition path sampling method, path integral path sampling (PIPS; Holdijk et al., 2024) which also trains a bias force by minimizing the KL divergence between path measures induced by the biased MD and the target path measure.
-For simplicity, we denote parameterizations for predicting force, potential, and atom-wise scaling as (F), (P) and (S), respectively.
-(a) Double-well energy landscapes
-(b) Distributions of the energy and $y$
-Figure 2: Visualization of potential energy landscapes and distributions of the double-well system.
-(2(a)) Visualization of the neural bias potential $b_{\theta}$ for TPS-DPS (P) in the double-well potential energy landscape.
-(2(b)) Distributions of the potential energy (left) and $y$ coordinate (right) of transition states from 1024 transition paths sampled by each method.
-4.1 Double-well system
-(a) UMD
-(b) SMD ($0.5$)
-(c) SMD ($1$)
-(d) TPS-DPS (F)
-(e) TPS-DPS (P)
-(f) Ground Truth
-Figure 3: 16 transition path sampled from each method on the potential energy landscape of the double-well system.
-White circles indicate the meta-stable states and white stars indicate the saddle points.
-We sample paths from the left meta-stable state to the right.
-SMD ($k$) denotes steered MD with the force constant $k$.
-All paths are generated by MD simulation at 300K.
-SMD ($k$) denotes steered MD with the force constant $k$.
-We begin by evaluating our method on a two-dimensional synthetic system, i.e., a double-well potential at 1200K.
-This system has two global minima (white circles) representing the meta-stable states, and two reaction pathways via saddle points (white stars).
-We sample transition paths from the left meta-stable state $\bm{R}_{\mathcal{A}}$ to the right meta-stable states =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\bm{R}-\bm{R}_{\mathcal{B}}\rVert<0.5\}$, using over-damped Langevin dynamics.
-We collect ground truth path ensembles by rejection sampling which proposes paths sampled from the unbiased MD simulations and accepts if the final states are in the target meta-stable states $\mathcal{B}$.
-We provide more details on the system in Section B.4.
-In Table 1, TPS-DPS shows superior performance for the double-well potential regardless of the bias force design.
-In Figure 2, the neural bias potential accelerates the transition by increasing the potential energy near the initial meta-stable state while decreasing the potential energy near the two energy barriers.
-Moreover, the distribution of energy and $y$ coordinates of the transition states from TPS-DPS is closest to the ground truth compared with other baselines, successfully capturing two reaction channels.
-In Figure 3, unbiased MD at 1200K fails to escape the initial state while SMD struggles to pass the saddle points.
-Our approach generates transition paths more similar to the ground truth than other methods.
-(a) UMD
-(b) SMD (10)
-(c) SMD (20)
-(d) PIPS (F)
-(e) TPS-DPS (F)
-(f) TPS-DPS (P)
-Figure 4: 16 sampled paths from each method on the Ramachandran plot of Alanine Dipeptide.
-White circles indicate meta-stable states and white stars indicate saddle points.
-We sample transition paths from the meta-stable state $C5$ (upper left) to $C7ax$ (lower right).
-SMD ($k$) denotes steered MD with the force constant $k$.
-Paths are generated by MD simulation at 1200K.
-Figure 5: Snapshots of transition path for Alanine Dipeptide (left) and Polyproline Helix (right).
-(a) UMD
-(b) SMD (10k)
-(c) SMD (15k)
-(d) PIPS (P)
-(e) TPS-DPS (P)
-(f) TPS-DPS (S)
-Figure 6: 16 sampled paths from each method projected to the top two TICA components.
-White circles indicate meta-stable states.
-We sample transition paths from the unfolded state (right) to the folded state (left).
-SMD ($k$) denotes steered MD with the force constant $k$.
-All paths are generated by MD simulation at 300K.
-(P) and (S) refer to predicting the bias potential and atom-wise scaling.
-4.2 Alanine Dipeptide conformation change
-We first consider Alanine Dipeptide for a real-world molecule consisting of two alanine residues, sampling transition paths from the $C5$ (upper left) to the $C7ax$ (lower right) as seen in Figure 4.
-The target meta-stable states are defined as =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$, where $\xi(\bm{R})=(\phi,\psi)$ is a well-known collective variable which consists of two backbone dihedral angles.
-Also, Alanine Dipeptide has two reaction channels between the $C5$ and $C7ax$ state passing through the saddle points marked as white stars.
-In Table 1 and Figure 4, unbiased MD at 300K fails to escape the initial state, while SMD with the collective variable $\xi$ generates transition paths with less probable transition states.
-PIPS generates transition paths of only one reaction channel, suffering from mode collapse.
-Compared to the baselines, our method successfully generates diverse transition paths that pass two reaction channels and show superior performance regardless of the bias force design.
-For qualitative analysis, we plot the snapshots of the transition path sample from TPS-DPS in Figure 5.
-4.3 Polyproline Helix isomerization
-Next, we consider a more complex molecule, Polyproline Helix consisting of three proline residues.
-We sample transition paths from the left-handed state (PPII) in the cis-configuration to the right-handed state (PPI) in the trans-configuration.
-For simplicity, we define the target meta-stable state based on handedness $H$ as =$\mathcal{B}=\{\bm{R}~{}|~{}H(\bm{R})>0\}$.
-For the formal definition of handedness, refer to Section B.4.
-As seen in Table 1, our models consistently outperform all baselines: unbiased MD, SMD, and PIPS.
-In Figure 7(a), a transition path sampled from TPS-DPS crosses the energy barrier around 0.4ps and changes handedness for 1.5ps.
-Additionally in Figure 5, we plot the snapshots of the sampled transition path.
-(a) Potential energy and CVs over time.
-(b) Visualization of hydrogen bond formation in Chignolin.
+TPS-DPS (S, Ours) 1.17 $\pm$ 0.66 59.380 -780.18 $\pm$ 216.93
+BBA
+UMD (400K) 10.03 $\pm$ 0.390 0.00 -
+UMD (1200K) 10.81 $\pm$ 1.050 0.00 -
+SMD (10k) 2.89 $\pm$ 0.32 0.00 -
+SMD (20k) 1.66 $\pm$ 0.30 26.560 -3104.95 $\pm$ 97.570
+PIPS (F) 9.84 $\pm$ 0.18 0.00 -
+PIPS (P) 9.09 $\pm$ 0.36 0.00 -
+TPS-DPS (F, Ours) 9.48 $\pm$ 0.18 0.00 -
+TPS-DPS (P, Ours) 3.89 $\pm$ 0.35 0.00 -
+TPS-DPS (S, Ours) 1.21 $\pm$ 0.09 84.380 -3801.68 $\pm$ 139.38
+Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
+Å % $\mathrm{kJmol}^{-1}$
+Trp-cage
+UMD (400K) 7.94 $\pm$ 0.65 0.00 -
+UMD (1200K) 8.27 $\pm$ 1.13 0.00 -
+SMD (10k) 1.68 $\pm$ 0.23 3.12 -312.54 $\pm$ 20.67
+SMD (20k) 1.20 $\pm$ 0.20 42.190 -226.40 $\pm$ 85.59
+PIPS (F) 7.47 $\pm$ 0.19 0.00 -
+PIPS (P) 6.07 $\pm$ 0.26 0.00 -
+TPS-DPS (F, Ours) 6.35 $\pm$ 0.31 0.00 -
+TPS-DPS (P, Ours) 3.15 $\pm$ 0.52 12.500 -512.97 $\pm$ 56.89
+TPS-DPS (S, Ours) 0.76 $\pm$ 0.12 81.250 0-317.61 $\pm$ 140.89
+BBL
+UMD (400K) 18.48 $\pm$ 0.630 0.00 -
+UMD (1200K) 18.90 $\pm$ 1.160 0.00 -
+SMD (10k) 3.67 $\pm$ 0.22 0.00 -
+SMD (20k) 2.97 $\pm$ 0.33 7.81 -1738.57 $\pm$ 386.81
+PIPS (F) 17.92 $\pm$ 0.290 0.00 -
+PIPS (P) 12.67 $\pm$ 0.310 0.00 -
+TPS-DPS (F, Ours) 10.15 $\pm$ 0.540 0.00 -
+TPS-DPS (P, Ours) 6.45 $\pm$ 0.26 0.00 -
+TPS-DPS (S, Ours) 1.60 $\pm$ 0.19 43.750 -3616.32 $\pm$ 213.66
+(a) Alanine Dipeptide
+(b) Chignolin
+(c) Visualization of Chignolin
 Figure 5: Qualitative evaluation on transition path sampled from TPS-DPS.
-(7(a)) Potential energies and collective variables (CVs) of Alanine dipeptide, Polyporline Helix, and Chignolin from top to bottom.
-For CVs, we plot the distance of the backbone dihedral angles between the current state and the target state for Alanine Dipeptide, the handedness of the backbone for Polyproline Helix, and the hydrogen bond distance for Chignolin.
-One can see that the sampled path successfully transitioned to the target meta-stable state.
-(7(b)) Visualization of hydrogen bond formation for Chignolin in a sampled path.
-We project a sampled path on the top two TICA components.
-We visualize the time steps indicated by the black circles, with each hydrogen bond highlighted in green and yellow.
-4.4 Chignolin folding
-Finally, we consider a challenging molecule, Chignolin, an artificial protein consisting of 10 amino acids (Honda et al., 2004), which folds into a $\beta$-hairpin structure by hydrogen bonds.
-We sample transition paths from the unfolded state (right white circle) to the folded state (left white circle) in Figure 6.
-We define the target meta-stable state =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$ where $\xi$ consists of the top two components of time-lagged independent component analysis (TICA; Pérez-Hernández et al., 2013).
+(5(a)) Potential energy and the two backbone dihedral angle distances between the current and target states.
+(5(b)) Potential energy and the two hydrogen bond distances between the current and the target state.
+(5(c)) Visualization of hydrogen bond formation in Chignolin.
+We highlight each hydrogen bond in green and yellow.
+4.3 Fast-folding Proteins
+Finally, for challenging molecules, we consider four fast-folding proteins (Lindorff-Larsen et al., 2011): Chignolin, Trp-cage, BBA, and BBL with 10, 20, 28, and 47 amino acids, respectively.
+Proteins fold into stable structures by forming networks of hydrogen bonds.
+We aim to sample protein folding processes as seen in Figure 7.
+We define the target meta-stable state as =$\mathcal{B}=\{\bm{R}~{}|~{}\lVert\xi(\bm{R})-\xi(\bm{R}_{\mathcal{B}})\rVert<0.75\}$ where $\xi$ consists of the top two components of time-lagged independent component analysis (TICA; Pérez-Hernández et al., 2013).
 We further describe TICA in Section B.4.
-[…]
+As shown in Table 2 and Figure 7, only TPS-DPS (S) successfully samples transition paths that pass probable transition states while TPS-DPS (F) and TPS-DPS (P) fail to hit the target meta-stable state due to the lack of meaningful training signal.
+While SMD hits the target meta-stable, its transition paths pass less probable transition states.
+UMD and PIPS fail to hit the target meta-stable state.
+In Figure 5, we validate the sampled paths using the potential energy and donor-acceptor distance of the two key hydrogen bonds in Chignolin folding, ASP3OD-THR6OG and ASP3N-THR8O used in (Yang et al., 2024).
+The sampled path of reduces the donor-acceptor distance below the threshold $3.5$Å.
+For 3D videos of transition paths of fast-folding proteins, we refer to project page.
 4.4 Ablation study
-Effectiveness of algorithmic components.
-First, we conduct ablation studies to verify the effectiveness of the five proposed components: log-variance loss, learnable control variate, replay buffer, simulated annealing, and maximum over RBF values for various path lengths.
-To be specific, we (1) replace our loss with the KL divergence, (2) replace the learnable control variate with the local control variate used in Nüsken & Richter (2021), (3) remove the replay buffer and use data only from the current policy, (4) use only one temperature $\lambda$, and (5) remove maximum operation over RBF values using only the final state.
-[…]
+We conduct ablation studies to verify the effectiveness of the six proposed components: log-variance loss, learnable control variate, replay buffer, simulated annealing, maximum over RBF values for various path lengths, and $SE(3)$ equivariance.
+To be specific, we (1) replace the log-variance divergence with the KL divergence, (2) replace the learnable control variate $w\in\mathbb{R}$ with the local control variate which is Monte-Carlo estimator used in Nüsken & Richter (2021), (3) remove the replay buffer and rely solely on data from the current policy, (4) use only one temperature $\lambda=300$K in sampling, (5) remove maximum operation over RBF kernel values using only the final state, (6) skip the alignment of neural network input states with the target position by the Kabsch algorithm.
 As seen in Figure 6, all the proposed components improve performance.
@@
 For the relaxed indicator function, maximum operation accelerates convergence and improves performance with frequent training signals from the subtrajectories.
+Leveraging the symmetry of the bias force with the Kabsch algorithm improves performance.
+We further compare with reverse KL divergence in Appendix E.
 Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
@@
 w/o various len 0.23 $\pm$ 0.11 75.00 29.49 $\pm$ 14.13
+w/o equivariance 0.34 $\pm$ 0.17 56.25 22.12 $\pm$ 16.96
 (a) Component-wise performance on Alanine Dipeptide.
-(b) Loss and RMSD curve over rollouts.
-Figure 9: Ablation studies on the components of TPS-DPS in the Alanine Dipeptide.
-(9(a)) Benchmark scores on Alanine Dipeptide.
-All metrics are averaged over 64 paths, and we highlight the best results in bold.
-ETS is computed for paths that hit the target meta-stable state.
-(9(b)) Loss and RMSD curves averaged over 8 different seeds.
-Comparison to steered MD with varying force constant $k$.
-The performance of SMD differs by the force constant $k$.
-As seen in Figure 8, one can see the trade-off between THP and ETS for various $k$ values.
-TPS-DPS outperforms SMD in THP and ETS without showing a trade-off between them.
+(b) Loss and RMSD curves over rollouts.
+Figure 6: Ablation studies on each proposed component of TPS-DPS (F) in Alanine Dipeptide. (6(a)) Benchmark scores on Alanine Dipeptide. (6(b)) Loss and RMSD curves averaged over 8 seeds.
 5 Conclusion
-In this work, we introduced a novel CV-free diffusion path sampler, called TPS-DPS, to amortize transition path sampling, using log-variance divergence with the learnable control variate, and the replay buffer with simulated annealing for off-policy training.
-We also propose new $SE(3)$ equivariant scale-based parameterization for bias force and relaxed indicator function for frequent training signals.
-Evaluations on synthetic double-well and real-molecule systems such as Alanine Dipeptide, Polyproline Helix, and Chignolin demonstrate superior accuracy and diversity of our model compared to both classical and ML approaches.
+In this work, we introduced a novel CV-free diffusion path sampler, called TPS-DPS, to amortize the cost of sampling transition paths.
+We propose the log-variance divergence with the learnable control variate and off-policy training with the replay buffer and simulated annealing.
+We propose a new scale-based equivariant parameterization of bias force and relaxed indicator function for reaching target meta-stable states and frequent training signals, particularly in large molecules.
+Evaluations on double-well, Alanine Dipeptide, and four fast-folding proteins demonstrate that TPS-DPS is superior in accuracy and diversity compared to non-ML and ML approaches.
 Limitation.
-While our algorithm successfully samples transition paths for small proteins, we did not evaluate our algorithm for even larger proteins due to the lack of computational budget.
-In addition, our work does not generalize across unseen pairs of meta-stable states or different molecular systems.
+While our experiments show promise, they are limited to small fast-folding proteins (up to 50 amino acids).
+The applicability of our method to real-world proteins with more than 500 amino acids remains unexplored.
+Additionally, integrating various libraries, such as PLUMED (plu, 2019) and DMFF (Wang et al., 2023), for MD with neural network bias force, has not yet been investigated.
+Our method does not generalize across unseen pairs of meta-stable states or different molecular systems.
 These points to an interesting venue for future research, which would be more appealing for practical applications in drug discovery or material design.
@@
 Following (Nüsken & Richter, 2021, Lemma A.1), we apply Girsanov’s Theorem to calculate the Radon-Nikodym derivative $\frac{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}{\mathrm{d}\mathbb{P}_{0}}$ as follows:
-$\frac{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\exp\left(\int^{T}_{0}(\bm{v}_{\theta}^{T}\Sigma^{-1})(\bm{X}_{t})\cdot\mathrm{d}\bm{X}_{t}-\int^{T}_{0}(\Sigma^{-1}\bm{u}\cdot\bm{v}_{\theta})(\bm{X}_{t})\mathrm{d}t-\frac{1}{2}\int^{T}_{0}\lVert\bm{v}_{\theta}(\bm{X}_{t})\rVert^{2}\mathrm{d}t\right)$ (14)
+$\frac{\mathrm{d}\mathbb{P}_{\bm{v}_{\theta}}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\exp\left(\int^{T}_{0}(\bm{v}_{\theta}^{T}\Sigma^{-1})(\bm{X}_{t})\cdot\mathrm{d}\bm{X}_{t}-\int^{T}_{0}(\Sigma^{-1}\bm{u}\cdot\bm{v}_{\theta})(\bm{X}_{t})\mathrm{d}t-\frac{1}{2}\int^{T}_{0}\lVert\bm{v}_{\theta}(\bm{X}_{t})\rVert^{2}\mathrm{d}t\right).$ (14)
 Since the state $\bm{X}_{t}$ follows the SDE $\mathrm{d}\bm{X}_{t}=(\bm{u}(\bm{X}_{t})+\Sigma\tilde{\bm{v}}(\bm{X}_{t}))\mathrm{d}t+\Sigma\mathrm{d}\bm{W}_{t}$.
@@
 […]
-A.3 Training algorithm
-Algorithm 1 Training
-1:  Initialize an empty replay buffer $\hat{\mathcal{D}}$, an policy $\bm{v}_{\theta}$, a scalar parameter $w$, the number of rollout $I$ and training per rollout $J$, and an annealing schedule $\lambda_{\text{start}}=\lambda_{1}>\cdots>\lambda_{I}=\lambda_{\text{end}}$.
-2:  for $i=1,\ldots,I$ do
-3:     Generate $M$ paths $\{\bm{x}_{0:L}^{(m)}\}^{M}_{m=1}$ from the biased MD simulations with $\bm{v}_{\theta}$ at temperature $\lambda_{i}$.
-4:     Update the replay buffer $\hat{\mathcal{D}}\leftarrow\hat{\mathcal{D}}\cup\{\bm{x}_{0:L}^{(m)}\}^{M}_{m=1}$.
-5:     for $j=1,\ldots,J$ do
-6:        Sample $K$ data $\{\bm{x}_{0:L}^{(k)}\}^{K}_{k=1}$ from $\hat{\mathcal{D}}$.
-7:        Update $\theta$ and $w$ with the gradient of $\frac{1}{K}\sum_{k=1}^{K}\left(\log\frac{p_{0}(\bm{x}^{(k)}_{0:L})1_{\mathcal{B}}(\bm{x}^{(k)}_{0:L})}{p_{\bm{v}_{\theta}}(\bm{x}^{(k)}_{0:L})}-w\right)^{2}$
-8:     end for
-9:  end for
+A.3 Proof of scale-based parameterization
+In this section, we prove that our scale-based parameterization of bias force strictly decreases the distance to the (aligned) target position for small step sizes, improving the ability to find informative paths in large molecules.
+Proposition 1.
+Consider the molecular state $\bm{R}_{t}$ at the $t$-th time step and the next state $\bm{R}_{t+\Delta t}^{\prime}=\bm{R}_{t}+\bm{b}(\bm{X}_{t})\Delta t/\bm{m}$ updated by step size $\Delta t$ and the bias force $\bm{b}(\bm{X}_{t})=\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))(\rho_{t}\cdot{\bm{R}}_{\mathcal{B}}-\bm{R}_{t})$.
+Then there always exists a small enough $\Delta t$ that strictly decreases the distance towards the target state $\bm{R}_{\mathcal{B}}$:
+$\lVert\rho^{\prime}_{t+\Delta t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert<\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert,$ (33)
+where $\rho^{\prime}_{t+\Delta t}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\rho\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert$ and we assume that there does not exist a rotation that exactly aligns the current molecular state to the target state, i.e., $\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert>0$.
+Proof.
+The proof consists of two steps.
+We first show the (strictly) positive correlation between the bias force and the direction from the $t$-th state $\bm{R}_{t}$ to the target state $\bm{R}_{\mathcal{B}}$.
+Next, we show that the positive correlation gaurantees a strict decrease in distance between the states, i.e., $\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert$, given that the distance was not already zero.
+Step 1: First, we show that the bias force (divided by atom-wise masses) is positively correlated with the direction to the target position, i.e., $(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})>0.$ This follows from:
+$(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$ $=(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})^{T}\frac{\text{diag}(\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t}))}{\bm{m}}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})$ (34)
+$=\sum_{i=1}^{3N}\left(\frac{\bm{s}_{i}}{\bm{m}_{i}}\right)\left(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\right)^{2}_{i}>0,$ (35)
+where $\bm{s}_{i}>0$ is the $i$-th element of $\bm{s}_{\theta}(\rho_{t}^{-1}\cdot\bm{X}_{t})$ and $\left(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\right)_{i}$ is the $i$-th element of the direction to the target position.
+Step 2: Next, we show that the positive correlation ensures distance reduction for a small enough step size.
+Consider the squared distance between the target position $\rho_{t}\cdot\bm{R}_{\mathcal{B}}$ and updated position $\bm{R}^{\prime}$ by bias force
+$\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert^{2}$
+$=\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-(\bm{R}_{t}+\bm{b}(\bm{X}_{t})\Delta t/\bm{m})\rVert^{2}$ (36)
+$=\lVert(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})-\bm{b}(\bm{X}_{t})\Delta t/\bm{m}\rVert^{2}$ (37)
+$=\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert^{2}-2\Delta t(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})+(\Delta t)^{2}\lVert\bm{b}(\bm{X}_{t})/\bm{m}\rVert^{2}.$ (38)
+Due to step 1, i.e., $(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})>0$, there exists a step size $\Delta t$ satisfying:
+$0<\Delta t<\frac{2(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t})}{\lVert\bm{b}(\bm{X}_{t})/\bm{m}\rVert^{2}}.$ (39)
+With this choice of $\Delta t$, multiplying $\Delta t\lVert\bm{b}(\bm{X}_{t})/\bm{m}\rVert^{2}$ leads to the following inequaliity:
+$(\Delta t)^{2}\lVert\bm{b}(\bm{X}_{t})/\bm{m}\rVert^{2}<2\Delta t(\bm{b}(\bm{X}_{t})/\bm{m})^{\top}(\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}).$ (40)
+By subtracting the right-hand side from both sides and adding $\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert^{2}$ to both sides, we have the following inequality:
+$\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert^{2}<\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert^{2}.$ (41)
+Taking the square root of both sides, we have the following inequality:
+$\lVert\rho^{\prime}_{t+\Delta t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert\leq\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert<\lVert\rho_{t}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{t}\rVert,$ (42)
+where the first inequality follows from the definition of $\rho^{\prime}_{t+\Delta t}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\rho\cdot\bm{R}_{\mathcal{B}}-\bm{R}^{\prime}_{t+\Delta t}\rVert$.
+This completes the proof.
+∎
 Appendix B Experiment details
 B.1 OpenMM configurations
-For real-world molecules, Alanine Dipeptide, Polyproline Helix, and Chingolin, we use the VVVR integrator (Sivak et al., 2014) with the step size $\Delta t=1~{}\mathrm{fs}$ and the friction term $\gamma=1~{}\mathrm{ps}^{-1}$.
-In the TPS-DPS training algorithm, we start simulations at a temperature $\lambda_{\text{start}}=600$K, and end at a temperature $\lambda_{\text{end}}=300$K.
-Other OpenMM configurations are shown in Table 2.
-Table 2: OpenMM configurations for three real-world molecular systems.
-Molecule Simulation time ($T$) Force field Solvent
-Alanine Dipeptide $1000$ amber99sbildn (Lindorff-Larsen et al., 2010) vaccum
-Polyproline Helix $10000$ amber/protein.ff14SBonlysc (Maier et al., 2015) implicit/gbn2
-Chignolin $5000$ amber/protein.ff14SBonlysc implicit/gbn2
+For real-world molecules, we use the VVVR integrator (Sivak et al., 2014) with the step size $\Delta t=1~{}\mathrm{fs}$ and the friction term $\gamma=1~{}\mathrm{ps}^{-1}$.
+In the TPS-DPS training algorithm, we simulate MD with $T=1$ps for Alanine Dipeptide and $T=5$ps for the fast-folding proteins.
+We start simulations at a temperature $\lambda_{\text{start}}=600$K, and end at a temperature $\lambda_{\text{end}}=300$K for Alanine Dipeptide and Chignolin and $\lambda_{\text{end}}=400$K for the others.
+We use the amber99sbildn (Lindorff-Larsen et al., 2010) force field for Alanine Dipeptide in vacuum and the ff14SBonlysc (Maier et al., 2015) force field for the fast-folding proteins with the gbn2 implicit solvation model (Nguyen et al., 2013).
 B.2 Model configurations
 We use a 3-layer MLP for the double-well system, and a 6-layer MLP for real-world molecules with ReLU activation functions for neural bias force, potential, and scale.
-To constrain the output of the neural bias scale parameterization to a positive value, we apply Softplus (Zheng et al., 2015) to the MLP output.
+To constrain the output of the neural bias scale parameterization to a positive value, we apply Softplus to the MLP output.
 As an input to the neural network, we concatenate the current position $(\bm{R}_{t})_{i}$ of the $i$-th atom with its distance to the target position $d_{i}=\lVert(\tilde{\bm{R}}_{\mathcal{B}})_{i}-(\bm{R}_{t})_{i}\rVert_{2}$.
@@
 (2024).
-For CVs of SMD, we use backbone dihedral angles $(\phi,\psi)$ for Alanine Dipeptide and RMSD for Polyproline Helix and Chignolin.
+For CVs of SMD, we use backbone dihedral angles $(\phi,\psi)$ for Alanine Dipeptide and RMSD for fast-folding proteins.
 Table 3: Model configurations of TPS-DPS.
@@
 Alanine Dipeptide 1000 16 16 1000 0.1
-Polyproline Helix 1000 16 4 200 0.2
-Chignolin 1000 16 4 200 0.5
+Chignolin 100 16 4 200 0.5
+Trpcage 100 16 4 100 0.5
+BBA 100 16 4 100 0.5
+BBL 100 16 2 100 0.5
 B.3 Evaluation metrics
@@
 B.4 System details
-Double-well potential and dynamics.
-Double-well system follows the overdamped Langevin dynamics defined as follows:
+Systhetic double-well system. Double-well system follows the overdamped Langevin dynamics defined as follows:
 $\mathrm{d}\bm{R}_{t}=\frac{-\nabla U(\bm{R}_{t})}{\bm{m}}\mathrm{d}t+\sqrt{\frac{2\gamma k_{B}\lambda}{\bm{m}}}\mathrm{d}\bm{W}_{t}\,.$ (45)
@@
 This potential has global minima and two saddle points, having two meta-stable states and two reaction channels.
-Handedness.
-For four points $A,B,C,D$, we define a handedness (Moradi et al., 2009) as follows:
-$H_{ABCD}=\frac{\overrightarrow{EF}\cdot(\overrightarrow{CD}\times\overrightarrow{AB})}{|\overrightarrow{EF}|\cdot|\overrightarrow{CD}|\cdot|\overrightarrow{AB}|}.$ (37)
-Here, $E$ and $F$ are the midpoints of the vectors $\overrightarrow{AB}$ and $\overrightarrow{CD}$, respectively.
-With the backbone atoms $X_{1},X_{2},...,X_{N}$ of a Polyproline Helix, we can define the handedness for $N$ atoms as $H=\sum_{i=1}^{N-3}H_{X_{i}X_{i+1}X_{i+2}X_{i+3}}$.
-In our experiments, we take the alpha carbon $X_{1}$, the carbonyl carbon $X_{2}$ of the first proline residue, the alpha carbon $X_{3}$ of the second residue and alpha carbon $X_{4}$ of third residues to compute the handedness.
-Time-lagged independent components (TICA).
 […]
-Reproducibility.
-We describe experiment details in Appendix B, including detailed simulation configuration and hyper-parameters. In the anonymous link, we provide the code for TPS-DPS.
+Reproducibility. We describe experiment details in Appendix B, including detailed simulation configuration and hyper-parameters. In the anonymous link, we provide the code for TPS-DPS.
+Appendix C Computational cost
+In this section, we analyze the time complexity of TPS-DPS and provide the number of energy evaluations and runtime in training and inference time for real molecules.
+The training and inference time complexity of TPS-DPS is $O(NMLJ)$ and $O(NML)$, respectively, where $N$ is the number of atoms, $M$ is the number of samples, $L$ is the number of MD steps, and $J$ is the number of rollouts.
+To be specific, training consists of biased MD simulations with $O(NML)$ time complexity.
+Given $M$, the total complexity of one biased MD step is $O(N)$.
+To justify it, we note that the biased MD step consists of three stages: (1) calculating bias force, (2) calculating OpenMM force field, and (3) integrating the biased MD.
+Given the number of layers, and hidden units, MLP for bias force requires $O(N)$ and the Kabsch algorithm for equivariance requires $O(N)$.
+Calculating force with cut-off and integrating MD with VVVR integrator also requires $O(N)$.
+To measure computational cost, we consider the number of energy evaluations and runtime per rollout in training and inference time.
+As shown in Table 4, the inference cost of TPS-DPS is proportional to UMD and SMD which have time complexity $O(NML)$.
+TPS-DPS requires less energy evaluations than PIPS in training since TPS-DPS finds transition paths faster than PIPS by utilizing the replay buffer and simulated annealing.
+Table 4: Cost comparison.
+EET and EEI refer to the number of energy evaluations by the OpenMM library in training and inference, respectively.
+RT and RI denote runtime (second) per rollout in training and inference on a single RTX A5000 GPU.
+MD simulations are conducted with $T=1$ps for Alanine Dipeptide, $T=5$ps for other systems, and $\Delta t=1$fs.
+The number of samples for each training rollout and inference is 16 and 64, respectively.
+Other configurations of TPS-DPS are provided in Section B.2 and those of PIPS can be found in (Holdijk et al., 2024)
+Molecule Method EET ($\downarrow$) EEI ($\downarrow$) RT ($\downarrow$) RI ($\downarrow$)
+Alanine Dipeptide UMD - 64K - 29.49
+SMD - 64K - 47.45
+PIPS (F) 240M 64K 44.22 71.05
+PIPS (P) 240M 64K 50.54 75.67
+TPS-DPS (F, Ours) 16M 64K 24.93 70.50
+TPS-DPS (P, Ours) 16M 64K 27.25 78.83
+TPS-DPS (S, Ours) 16M 64K 25.11 73.04
+Chignolin UMD - 320K - 224.23
+SMD - 320K - 380.37
+PIPS (F) 40M 320K 277.29 565.58
+PIPS (P) 40M 320K 317.08 622.87
+TPS-DPS (F, Ours) 8M 320K 209.29 562.90
+TPS-DPS (P, Ours) 8M 320K 224.36 623.63
+TPS-DPS (S, Ours) 8M 320K 215.18 581.26
+Trp-Cage UMD - 320K - 258.29
+SMD - 320K - 323.52
+PIPS (F) 40M 320K 360.55 652.59
+PIPS (P) 40M 320K 417.93 718.05
+TPS-DPS (F, Ours) 8M 320K 289.10 655.22
+TPS-DPS (P, Ours) 8M 320K 301.76 699.44
+TPS-DPS (S, Ours) 8M 320K 293.51 673.00
+BBA UMD - 320K - 395.12
+SMD - 320K - 551.04
+PIPS (F) 40M 320K 506.57 1080.62
+PIPS (P) 40M 320K 574.07 1117.74
+TPS-DPS (F, Ours) 8M 320K 422.23 1042.81
+TPS-DPS (P, Ours) 8M 320K 430.24 1091.97
+TPS-DPS (S, Ours) 8M 320K 426.48 1068.68
+BBL UMD - 320K - 673.55
+SMD - 320K - 853.77
+PIPS (F) 40M 320K 622.22 1558.53
+PIPS (P) 40M 320K 660.02 1612.78
+TPS-DPS (F, Ours) 8M 320K 560.95 1520.05
+TPS-DPS (P, Ours) 8M 320K 572.77 1607.62
+TPS-DPS (S, Ours) 8M 320K 563.45 1553.89
+Appendix D Visualization of sampled paths of fast folding proteins
+(a) UMD
+(b) SMD (20k)
+(c) PIPS (P)
+(d) TPS-DPS (F)
+(e) TPS-DPS (P)
+(f) TPS-DPS (S)
+Figure 7: Visualization of sampled paths of four fast-folding proteins on free energy landscapes for top two TICA components.
+We aim to sample folding processes for the four fast folding proteins: Chignolin, Trpcage, BBA, and BBL (from top to bottom rows).
+Appendix E Comparision with reverse KL divergence
+Table 5: Benchmark scores of reverse KL divergence and TPS-DPS on Alanine Dipeptide system.
+Metrics are averaged over 64 paths.
+TPS-DPS consistently outperforms reverse KL divergence on all metrics regardless of predicting bias force or potential.
+Method RMSD ($\downarrow$) THP ($\uparrow$) ETS ($\downarrow$)
+Å % $\mathrm{kJmol}^{-1}$
+Reverse KL (F) 0.43 $\pm$ 0.34 53.12 27.88 $\pm$ 14.38
+Reverse KL (P) 0.58 $\pm$ 0.34 48.43 21.61 $\pm$ 11.76
+TPS-DPS (F, Ours) 0.16 $\pm$ 0.06 92.19 19.82 $\pm$ 15.88
+TPS-DPS (P, Ours) 0.16 $\pm$ 0.10 87.50 18.37 $\pm$ 10.86
+(a) Reverse KL (F)
+(b) Reverse KL (P)
+(c) TPS-DPS (F)
+(d) TPS-DPS (P)
+Figure 8: Visualization of sampled paths of Alanine Dipeptide on the Ramachandran plot.
+The reverse KL divergence struggles to find diverse reaction channels suffering from mode collapse issues while the log-variance divergence of our method can capture two reaction channels and reach the target states better.